A Complete Bibliography of Philosophy of Science: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

08 November 2014 Version 1.01

Title word cross-reference

1700 [49]. 17th [303]. 18th [303]. 1900 [205]. 1912 [175]. 1933 [91]. 1973 [520]. 1974 [240].

2 [225].

60th [672].

90th [376].

Absolute [218]. Abstract [304]. Abusing [317]. Abyss [430]. Acceleration [14]. Acceptance [297]. Accepting [295]. Account [199]. Accuracy [475]. Achinstein [626, 334, 134]. Ackermann [484]. Acock [332]. Acting [486]. Action [486]. Addendum [684]. Additivity [251]. Adequacy [209]. Adequate [585, 194]. Adolf [440]. Adrienne [91].

Advancement [687]. Aeltere [412]. Affairs [138]. After [102, 156, 91]. After-Trial [102]. Against [344, 483, 515, 4, 317]. Agassi [241]. Agazzi [337]. Age [215, 332]. Ages [604, 560]. Agree [558]. Aim [283]. Aims [533, 481]. Air [516]. Air-Pump [516]. Akademie [412]. Alexander [487, 202, 138]. Algebraic [70]. Allan [605]. Alleged [94]. Allen [320, 71]. Almeder [433]. Alternative [479]. Alvin [624]. Alwyn [518]. Ambiguities [83, 168]. Ambition [497]. Amelie [139]. American [19]. among [211, 499]. Ampliative [679]. Analogy [299]. Analysis [346, 213, 426, 233, 329, 284, 352]. Analytic [450]. Analyzing [512, 422]. Ancient [337]. Andrew [427]. Anemia [619]. Animalium [17]. Anomalies [135]. Anomalous [135]. Antike [412]. Antiquity [604]. Apel [451]. Apologia [279]. Applications [537, 220]. Applying [117]. Approach [42, 311, 677, 70, 484]. Approximate [529, 133]. Approximation [252]. Approximative [365]. Archibald [407]. Archival [91]. Argument [58, 443, 556, 4, 477, 360, 527, 683, 538]. Arguments [63]. Aristoteles [412]. Aristotelian [165]. Aristotle [17]. Armstrong [495, 452]. Aronson [409]. Arpad [281]. Arthur [261, 34, 590]. Artifact [161]. Ascription [680]. Ash [91]. Aspects [562]. Association [291]. Assumptions [45]. Asymmetries [326]. Atomic [408, 148]. Atomism [117]. Automation [569]. Averages [39]. Averaging [668]. Avoiding [251]. Axiomatics [617, 337]. Axiomatization [437]. Axioms [620].

B [20, 258, 306, 320]. Back [21, 37, 52, 74, 92, 109, 126, 140, 159, 179, 191, 206, 226, 247, 263, 286, 307, 322, 338, 355, 378, 396, 415, 434, 453, 470, 488, 504, 525, 546, 563, 576, 591, 608, 628, 642, 658, 674, 693, 708, 87]. Backwards [94, 95]. Baigrie [619]. Baker [627]. Band [412]. Barbara [394]. Barker [134]. Barone [575]. Barry [353]. Bartley [283]. Barwise [704]. Bas [173]. Based [14]. Bateson [157]. Bayesianism [647]. Be [595, 150]. Bealer [333]. Beatty [682]. Beauregard [705]. Becoming [524, 318]. Beginning [581]. Beginnings [281]. Behalf [273]. Behavior [217, 690]. Behaviorism [213, 353]. Behavioristic [417]. Being [237, 318]. Belief [41, 445, 262, 382, 23, 515]. Bell [447, 568, 477, 613]. Bell-Epr [447]. Benacerraf [413]. Bernhard [500]. Bertrand [426]. Best [59, 102]. Better [273]. Between [5, 312, 123, 224, 251]. Beyond [84, 380, 573, 245]. Biography [16]. Biological [405, 195, 487, 313, 473]. Biology [55, 502, 325]. Biomedical [393]. Biophilosophical [500]. Birth [483]. Birthday [376, 519, 672]. Black [647, 548, 175]. Block [225]. Body [94, 154, 175, 548]. Boër [641]. Bohr [485]. Book [155, 222, 58, 90, 260, 524, 405, 336, 705, 656, 244, 486, 433, 281, 688, 370, 375, 391, 205, 35, 189, 259, 468, 276, 604, 623, 157, 372, 407, 483, 541, 280, 689, 588, 502, 20, 431, 427, 71, 640, 690, 707, 304, 692, 242, 282, 487, 440, 671, 221, 428, 158, 139, 108, 334, 572, 501, 480, 406, 408, 624, 125, 258, 451, 687, 156, 390, 625, 574, 18, 49, 394, 376, 448, 704, 605, 47, 173, 409, 32, 540, 283, 429, 518, 91, 333, 657, 371, 15, 626, 318, 542, 573, 220, 302, 352, 19, 153]. Book [36, 414, 305, 335, 589, 124, 543, 520, 450, 430, 306, 393, 466, 123, 17, 389, 432, 410, 639, 69, 176, 152, 245, 521, 627, 154, 485, 388, 350, 224, 515, 241, 240, 500, 70, 561, 203, 16, 223, 73, 377, 559, 523, 121, 277, 72, 261, 374, 392, 519, 672, 562, 175, 122, 204, 387, 262, 351, 243, 89, 522, 50, 33, 225, 319, 354, 284, 497, 34, 138, 317, 498, 320, 655, 251, 413, 279, 411, 412, 575, 278, 68, 137, 353, 174, 449, 202, 369, 686, 190, 706, 303, 641, 654, 481, 337, 496, 590, 691]. Book [48, 560, 373, 452, 503, 499, 465, 516, 467, 517, 285, 484, 482, 544]. Books [178, 246, 321, 395, 469, 545, 606, 673]. Bootstrap [557, 599, 270, 442]. Bourgeois [36]. Box [647]. Boyle [516]. Brain [631, 689]. Brains [148]. Brainstorms [32]. Branch [382]. Brand [486]. Brian [108, 49, 369]. Brings [554]. British [19]. Broglie [376]. Brown [692, 48, 655]. Brownian [615]. Bruce [502]. Buchwald [562]. Building [55]. Burdens [687]. Burks [34]. Buroker [242].

C [232, 656, 688, 468, 574, 173, 32, 70, 284, 174, 691]. Calculus [335]. Calibration [385]. Can [534]. Cantor [205, 18]. Cantorian [623]. Capacity [680]. Capital [203]. Caplan [88]. Carl [672]. Cartwright [531, 406]. Case [607, 695, 614, 152, 515, 549, 475, 317]. Categories [405, 117]. Causal [601, 439, 572, 211, 565, 699, 472, 527, 681, 423, 648, 80, 108]. Causality [711, 210, 514, 132, 399, 219, 364, 669, 172]. Causation [541, 269, 534, 699, 424, 509, 278, 145, 366]. Cause [34, 161, 310, 555, 664, 632]. Causes [458]. Causey [97, 68]. Centuries [303]. Century [90, 371, 72, 158, 562]. Chance [309, 262, 682, 350, 34]. Change [41, 26, 284, 678, 337]. Chapman [352]. Character [290]. Charity [231]. Charles [433, 496]. Chattopadhyaya [354]. Chauvinism [493]. Chemical [373]. Cherniak [654]. Choice [26]. Chomsky [686, 89, 686, 224]. Christopher [671, 654]. Churchland [594, 448, 95]. Churchlands [386]. Circularity [704]. Claim [421]. Claims [186]. Clark [121, 33, 346]. Clarke [523]. Class [87, 250]. Classical [80]. Claw [137]. Clock [218]. Cochrane [430]. Coefficients [239]. Coenzymes [312]. Cognition [540, 624]. Cognitive [540, 403, 515, 663, 551]. Cohen [220, 241, 377, 392, 411, 511]. Coherence [385, 507]. Collapse [568]. Collections [649]. Collegio [522]. Collier [366]. Colodny [561]. Commensurability [652, 651]. Comment [7, 441]. Commentary [17, 156]. Comments [571, 97, 315]. Commitment [48]. Common [195, 458, 632, 310, 555, 664, 528]. Communication [296]. Community [373]. Comparing [632]. Compatible [638]. Competing [607, 549]. Complementarity [361, 485]. Complexity [318, 320]. Components [614]. Computation [540]. Computational [579, 386]. Concept [611, 7, 47, 472, 193, 517, 333]. Conception [585, 475]. Conceptions [205]. Concepts [558, 293, 261, 422]. Conceptual [284]. concerning [418]. Condition [212]. Conditional [184, 249, 212]. Conditionalization [13, 129]. Conditionals [262, 367]. Conditioning [57, 420, 111]. Confidence [101, 102]. Confirmation [557, 103, 600, 702, 358, 133, 603]. Confirmational [129]. Conflict [457, 251].

Confutation [78]. Connections [312]. Consciousness [448]. Consensus [661, 479]. Consequence [106]. Consequences [423]. Consilience [383]. Consolidation [70]. Constructing [428, 427]. Constructions [492]. Contemporary [426, 448, 70, 15]. Content [9, 671]. Context [426]. Continuity [678, 331]. Continuum [604]. Contrast [587]. Contributions [15]. Controversy [83, 703, 499]. Convention [476, 375]. Conventionalist [46]. Conventionality [349]. Convergence [660, 597]. Convergent [198, 78]. Copenhagen [404]. Copernican [475]. Correction [607, 616]. Corrections [51, 177]. Correlation [419]. Correlations [637, 249, 424]. Corroborated [167]. Corroboration [278]. Cosmology [676, 589]. Costa [705]. Countable [649]. Counterfactual [447, 145]. Counterfactuals [399, 509]. Counterthemes [371]. Craven [17]. Creation [604, 498, 676]. Creationism [317]. Credal [350]. Crisis [117]. Crispin [124]. Criterion [107, 65]. Critical [275, 685, 603]. Criticism [172]. Critique [699, 46, 440]. Critiques [645]. Crosland [36]. Crossed [586]. Crossroads [502]. Cummins [346, 503]. Curd [475]. Current [117, 380]. Currie [542].

D [280, 574, 376, 450, 50, 138, 337, 452]. Danger [117]. Daniel [32, 203, 691]. Dark [430]. Darwin [214, 527, 690]. Darwinian [137, 598]. Darwinism [371, 138]. Data [233, 529, 484]. Dauben [18]. David [155, 502, 245, 521, 122, 495]. Davis [50]. Dawkins [319]. Debate [168, 224, 380, 494, 481]. Debiprasad [354]. Debra [625]. Decision [458, 402, 266, 262, 423, 648]. Decisive [360]. Deduction [541]. Deductive [281, 365, 683]. Deductive-Nomological [365]. Deductivism [51, 31, 201, 554]. Defence [483]. Defense [524, 269, 125, 443, 198, 101, 665, 216]. Defined [422]. Defining [677]. Definiteness [368, 145, 267]. Definition [66, 170]. del [279]. Demopoulos [684]. Denbigh [482, 482]. Dennett [32, 691]. Dependent [269, 211, 534, 366]. Depew [502]. Descarte [523]. Descriptions [217]. Descriptive [382, 144]. Design [55, 571, 384]. Designators [272, 537]. Desmond [523]. Determinational [681]. Determining [703]. Determinism [9, 132, 293, 10, 146, 681, 461, 685]. Deterministic [472]. Devonian [499]. Dialectical [484]. Did [366, 464]. Dieks [10]. Dimensions [371]. Diner [376]. Direct [271, 129]. Directed [217]. Directions [565]. Disagreement [479]. Disappearance [640]. Discontinuity [175]. Discovery [552, 312, 679, 553, 152, 381, 163, 380, 30]. Discussion [222]. Disorder [130]. Disposition [422]. Dispositions [422]. Disproof [701]. Distant [67]. Distributions [249, 348, 613]. Disunity [661]. Divorce [380]. Do [162, 613]. Does [599, 616, 404, 442]. Domains [529]. Dominance [661]. Doppelt [586]. Double [464]. Drake [16]. Dretske [176]. Dualism [376]. Dudley [387]. Dump [201]. Dupré [514]. Dutch [58, 251]. Dwight [50]. Dynamic [507].

E. [232]. Early [604, 291, 158, 303]. Earman [33, 603, 685]. Earth [430].

Ecological [320]. Economics [330, 112, 446, 441, 203]. Economist [330]. Economizing [97]. Economy [463]. Ed [603]. Edelson [466]. Edward [408]. Eells [711]. Effects [663, 239]. Efficacy [527]. Eighteenth [158]. Einstein [428, 589, 590, 360, 476]. Elbow [691]. Electricity [303]. Electromagnetic [562]. Elementary [300]. Elements [281]. Eliot [72]. Ellery [711]. Elliott [431, 344]. Ellsworth [15]. Elucidations [156]. Embeddings [602]. Embryology [656]. Emergence [690, 293]. Emerton [501]. Emigration [91]. Emotions [139]. Empirical [9, 209, 347]. Empirically [194]. Empiricism [182, 401]. Empiricist [155, 153]. Enc [106]. End [3]. Energy [589]. Enterprise [350]. Entropy [490, 528, 482]. Entwicklung [520]. Enzyklopadie [244]. Epistemic [601]. Epistemological [232]. Epistemology [114, 672, 256, 30, 347, 294, 588, 624, 377, 449]. Epr [447, 568, 637, 404]. Equilibria [253]. Ergodic [39]. Erlichson [171]. Ernan [47, 498]. Ernest [222]. Error [273]. Errors [195]. Essay [671, 704, 410, 350, 203, 285]. Essays [483, 374, 519, 672, 259, 17]. Essentialism [630, 40]. Essler [672]. Established [268]. Etchemendy [704]. Ether [205]. Ethics [382]. Ethnocentrism [166]. Ethnoscience [166]. Etiologies [106]. Euclid [281]. Euclidean [229]. Evaluation [383, 393]. Evelyn [394]. Event [167]. Everyone [534, 366]. Evett [261]. Evidence [512, 327, 667, 466, 95, 511, 664, 121, 277]. Evidential [255]. Evidentialist [417]. Evolution [96, 678, 40, 315, 253, 115, 500, 502, 498]. Evolutionary [55, 431, 690, 371]. Existence [410, 613]. Expectancy [64]. Expectations [56, 298]. Expected [633]. Experience [95]. Experiences [94, 168]. Experiment [626, 404, 605]. Experimental [171, 95, 72, 79, 516]. Experimentally [703]. Experimenter [64]. Experiments [136, 1, 571, 462, 384]. Explain [616]. Explaining [195, 603, 139]. Explanation [66, 62, 45, 572, 25, 408, 451, 59, 167, 565, 399, 491, 393, 341, 365, 584, 473, 278, 310, 646, 508, 334, 503, 541]. Explanations [42, 80]. Explanatory [128, 326]. Explicate [529]. Explicating [635]. Explicativity [297]. Exploratory [233, 401]. Expository [285]. Extended [319]. Extensive [29]. Extragalactic [695].

F [20, 258, 376, 320, 285]. Factors [83]. Facts [44]. Failure [150]. Fallacy [348]. False [298]. Faraday [428]. Fargue [376]. Farrell [306]. Faust [521]. Favoring [360]. Feasible [100]. Feedback [96]. Feeling [394]. Feigl [240]. Fernandes [688]. Feshbach [519]. Fetzer [278]. Feyerabend [615]. Fiducial [81]. Field [349, 125, 13]. Fifth [519]. Fine [419, 590]. Finite [251]. First [72, 347]. Fitnes [316]. Fitness [170, 478]. Five [332]. Fixing [54]. Flaw [134]. Flow [622, 176]. Focus [431]. Fodor [593]. Folk [515]. Folse [485]. Force [165]. Forces [636]. Forensic [512]. Form [501]. Formation [373]. Formulations [472]. Fortuitous [252]. Foss [539]. Foundation [540]. Foundations [405, 440, 124, 190, 421, 688, 33, 465]. Four [62]. Fox [394]. FPL [653]. Fraassen [539, 295, 173]. Framework [181, 485]. Francesco [575]. Francia [260]. Frankel [172]. Franklin [605]. Frantz [20]. Freaks [147]. Fred [541, 176]. Free [691]. Frequency [269, 534, 366]. Frequency-Dependent [269, 534]. Friedman [684, 465]. Front [22, 38, 53, 76, 93, 110, 127, 142, 160, 180, 192, 208, 227, 248, 264, 288, 308, 323, 339, 357, 379, 397, 416, 436, 454, 471, 489, 506, 526, 547, 564, 578, 592, 609, 629, 644, 659, 675, 694, 710]. Fuhrman [15]. Function [631]. Functional [346, 495, 11, 200, 472, 151]. Functionalism [362]. Functions [327, 665]. Future [456]. Fuzzy [2].

G [260, 205, 376, 542, 627, 561, 672, 641, 482]. Gale [73]. Galileo [223, 522, 584, 337, 331, 16, 522]. Game [590]. Garcia [543]. Garden [429]. Garland [71]. Garrett [405]. Gaston [19]. Gauge [463, 300]. Gay [36]. Gay-Lussac [36]. Gene [319]. General [425]. Genetic [114]. Genetics [200, 311, 157]. Genic [239, 161, 274]. Gentlemanly [499]. Geometrical [131]. Geometry [437, 349]. Georg [18]. George [333, 450, 73]. Georgia [451]. Gerald [336, 687]. German [91, 373]. German-Speaking [91]. Geschichte [412]. Get [218]. Gibbins [188, 706]. Gibbs [668, 39]. Gibson [285]. Giedymin [375]. Giere [269, 424, 366]. Gillies [556]. Gino [518]. Given [212]. Glenn [262]. Global [663]. Glymour [33, 557, 255, 103, 121]. Goal [217]. Goal-Directed [217]. Goldman [624]. Gonzalo [372]. Goodman [8, 134]. Grammar [627]. Grand [463, 300]. Grattan [335]. Grattan-Guinness [335]. Gravitational [695]. Great [275, 499]. Greek [281]. Green [560]. Green-Pedersen [560]. Grene [625, 371]. Group [566, 476]. Gruender [337]. Grunbaum [86, 440]. Grundriss [412]. Guide [91]. Guinness [335].

H [280, 502, 125, 258, 91, 519, 320, 278]. Hacker [627]. Hacking [389]. Hahn [391]. Half [90]. Hall [123]. Hallett [623, 477]. Hamilton [243]. Hankins [243]. Hannaway [626]. Hans [153]. Happen [366]. Harman [282]. Harold [692, 48, 620]. Harper [262]. Harre [172, 199, 422]. Hartry [125]. Harvey [106]. Hausman [203, 441]. Hawrylchak [91]. Healey [221]. Health [7, 6]. Hearst [72]. Heelan [370]. Hegel [411]. Heilbron [303]. Heister [652]. Hempel [672, 107]. Hennig [50]. Henry [259, 485, 449]. Herbert [240]. Heritage [522]. Hermeneutic [228]. Hesse [276, 220]. Hidden [568]. Hierarchical [457, 633]. Hierarchy [320]. Hilary [588, 672, 351, 413]. Hintikka [337]. Hints [249]. Histoire [543]. Historical [45, 426, 28, 475, 216]. Historiography [544]. History [232, 656, 205, 483, 427, 430, 84, 360, 354, 544, 335, 351]. Hobbes [131, 516]. Hodge [205]. Holism [663]. Holland [544]. Holographic [457]. Holton [336, 687]. Homunculus [27]. Honig [573]. Honor [519]. Honour [672]. Hooke [430]. Hooker [70]. Hope [554]. Hopeless [51, 31]. Horder [656]. Horwich [567, 277]. Hubert [407]. Hufbauer [373]. Hugh [353]. Hull [346, 213]. Human [542, 497, 138, 353, 24]. Humanism [49]. Hunt [71]. Husain [390]. Husserl [550]. Husserlian [530]. Hypotheses [632, 297]. Hypothesis [460, 626, 204, 466]. Hypothetical [97]. Hypothetico [51, 31, 201, 554]. Hypothetico-Deductivism [51, 31, 201, 554].

Ian [281, 389]. Idealism [242]. Idealistic [410]. Idealized [529]. Identifiable [649]. Identifying [273]. Identities [97]. Identity [197, 119, 612, 610, 235]. Ideological [645]. Ideology [325]. If [366]. Ifs [262]. II [70, 174]. Ilkka [432]. Illusion [117]. Ilya [318]. Image [539, 295, 173]. Imagery [494]. Images [284]. Implications [500]. Importance [475]. Imprecision [400]. Incommensurability [580]. Incompatible [359]. Incomplete [632, 482]. Incongruence [242]. Inconsistency [548]. Inconsistent [254]. Indefinite [570]. Independence [249]. Independent [1, 171, 664]. Indeterminacy [400, 328]. Indeterminism [219, 509]. India [354]. Indiscernibles [197, 119, 235]. Individualism [491]. Individuality [617]. Individuals [617, 421]. Induction [59, 556, 701, 163, 134, 574, 305]. Inductions [383]. Inductive [220, 511, 174]. Inequalities [613]. Inertia [99, 296]. Inference [679, 238, 271, 81, 129, 653, 449]. Inferences [100]. Infernal [82]. Infinite [662, 18]. Infinity [621, 510]. Information [75, 141, 207, 287, 356, 435, 505, 577, 643, 709, 104, 528, 508, 176]. Information-Theoretic [508]. Innovation [661]. Inorganic [500]. Inquiries [240]. Inquiry [372, 533, 336, 480]. Instrumental [666]. Instrumentation [228]. Instruments [484]. Intending [486]. Intensional [61, 657]. Intensity [464]. Intentional [680]. Intentionality [666]. Interaction [439]. Interest [426]. Interference [184]. Interfield [312]. Interpenetration [325]. Interpretation [316, 181, 404]. Interpretive [17]. Intertheoretic [456]. Intervals [101, 102]. Intervening [232, 389]. Intrinsic [348]. Introduction [448]. Introductory [335, 389]. Introspection [640]. Invariance [12, 477]. Invention [236, 163]. Investigation [260]. Irreversibility [181]. Isaac [350]. Isotropy [703, 476]. Issues [15, 342].

J [656, 205, 502, 690, 304, 428, 156, 589, 485, 559, 33, 655, 303, 337, 560, 499, 482, 544]. J.L [90]. Jackson [67]. James [278]. Janis [86]. Jardine [483]. Jean [543, 224]. Jed [562]. Jeffrey [13, 111, 174]. Jeffreys [620]. Jerrold [409]. Jerry [19, 593]. Jerzy [375]. Jill [242]. John [407, 704, 91, 33, 467, 484, 603, 514, 112, 685]. Joint [613]. Jon [704]. Jonathan [220]. Joseph [18, 241, 284]. Judgment [570, 400]. Jurgen [244]. Justification [601, 236, 30, 24]. Justificational [460]. Justifying [238].

Kant [688, 242, 296]. Karl [451, 283, 275, 373]. Karl-Otto [451]. Katz [304]. Keller [394]. Kennedy [136, 1]. Kennedy-Thorndike [136]. Kepler [483]. Kinds [147, 633]. Kinematics [58, 41, 667, 507, 23]. Kinetic [552, 149]. King [158]. Kitcher [388, 497, 315, 317]. Kitts [88]. Knowing [641]. Knowledge [688, 372, 350, 143, 278, 686, 499, 212, 388, 387, 482, 176, 686]. Kordig [257]. Kornblith [588]. Krajewski [374]. Kripke [302]. Kuhn [175]. Kuhnian [117]. Kyburg [81, 272, 449, 532].

L [524, 688, 258, 409, 220, 414, 392, 262, 243, 68, 303]. Lacey [353]. Language [686, 528, 686, 302, 405, 304, 224]. Larry [689, 393, 204, 481]. Last [562, 464]. Laszlo [519]. Later [550, 90]. Laudan [460, 393, 479, 163, 392, 204, 481]. Law [296, 511, 452, 517]. Lawhood [635]. Lawlikeness [3]. Lawrence [707]. Laws [66, 28, 495, 406, 113, 358, 216]. Laymon [615]. Learning [224, 417]. Lectures [686]. Leeds [82]. Leibniz [123, 655]. Lennon [156, 156]. Lensing [695]. Lester [158]. Let [201]. Levi [570, 271, 350, 533, 532]. Leviathan [516]. Levin [154]. Lewis [699]. Liar [704]. Libet [120]. Lie [406]. Life [394, 516]. Light [703, 402, 589]. Likelihood [555, 597]. Limitation [623]. Limited [46]. Limits [372, 521, 706, 336]. Link [312]. Loar [369]. Local [663, 145]. Lochak [376]. Logic [177, 182, 184, 702, 85, 429, 118, 188, 553, 163, 477, 183, 382, 174, 185, 267, 144, 554, 35, 220, 706]. Logical [153, 548]. Logico [70, 575]. Logico-Algebraic [70]. Lopes [90]. Lorentz [360, 476]. Louis [376]. Luce [43]. Lüders [185]. Lussac [36]. Lycan [641]. Lydia [430]. Lyons [640].

M [90, 281, 205, 282, 408, 156, 448, 91, 573, 627, 203, 452]. Machine [82, 63, 551]. MacKinnon [408]. Macro [461]. Macro- [461]. Macroeconomics [342]. Macroscopic [181]. Madden [422, 199]. Magnitude [705]. Malament [668]. Malebranche [156]. Man [71]. Managua [686]. Mandrou [49]. Marcello [279]. Margenau [259]. Marjorie [625, 371]. Marr [579]. Marshall [466]. Martha [17]. Martin [499]. Marx [241, 377, 411]. Mary [276, 220]. Massimo [224]. Materialism [164]. Materials [91]. Mathematical [388, 612]. Mathematics [281, 18, 265, 281, 657, 124, 413, 496]. Matter [21, 37, 52, 74, 22, 38, 53, 76, 92, 109, 126, 140, 93, 110, 127, 142, 159, 179, 191, 206, 160, 180, 192, 208, 226, 247, 263, 286, 227, 248, 264, 288, 307, 322, 338, 355, 308, 323, 339, 357, 378, 396, 415, 434, 379, 397, 416, 436, 453, 470, 488, 504, 454, 471, 489, 506, 525, 546, 563, 576, 526, 547, 564, 578, 591, 608, 628, 642, 592, 609, 629, 644, 658, 674, 693, 708, 659, 675, 694, 710, 426, 47, 448]. Maurice [36]. Maximal [107]. Maximization [633]. Maxwell [562, 552]. McCauley [98]. McClintock [394]. McMullin [47, 498]. Meaning [428, 257, 122, 369]. Meaningfulness [43]. Measurement [611, 29, 697, 407]. Mechanical [569]. Mechanics [14, 90, 611, 296, 70, 360, 368, 429, 190]. Medicine [158, 6, 393]. Mellor [280]. Memory [689]. Mental [130]. Merwe [518]. Metaphor [343]. Metaphysical [567, 474]. Metaphysics [189, 410, 639, 282, 154]. Metasociological [15]. Method [661, 390, 559]. Methodological [438, 403, 386, 646]. Methodology [377, 337, 672]. Metodo [279]. Metrical [582]. Michael [623, 15, 154, 137, 465]. Michelson [136, 1]. Micro [293, 461]. Micro-Determinism [293, 461]. Microeconomics [342]. Microphysics [562]. Middle [604, 560]. Mill [112, 441]. Miller [443, 15, 701, 683]. Millikan [405]. Mind [94, 690, 700, 154, 448, 393, 369]. Mind-Body [94, 154]. Minimal [654]. Minkowski [229, 437]. Minnesota [33]. Minogue [61]. Minutes [332].

Mirror [69]. Mittelstaedt [35]. Mittelstrass [244]. Mode [253]. Model [55, 479, 4, 653, 508]. Model-Theoretic [4]. Modeling [457, 565]. Models [55, 417, 566, 551, 419]. Modern [189, 47, 626, 561, 303, 363, 253, 583, 429]. Modest [425]. Modification [297]. Moment [344]. Monism [344]. Morgan [71]. Morison [336]. Moritz [258]. Morley [136, 1]. Motion [615, 296]. Motu [17]. Muddle [275]. Mueller [281]. Mulder [258]. Multifarious [199]. Munévar [372]. Musgrave [542]. Must [558, 595]. My [464]. Myles [486]. Mystery [637].

N [205, 483, 33]. Nadeau [189]. Nagel [66]. Nails [625]. Nancy [428, 406]. Narrow [362]. Nathan [524]. Nations [430]. Natural [309, 282, 147, 596, 228, 550, 113, 389, 214, 374, 519, 616, 256, 655, 682, 199]. Naturalism [155]. Naturalized [486, 398, 256, 347, 294]. Naturalizing [30, 588]. Nature [189, 372, 431, 147, 334, 113, 313, 114, 214, 388, 686, 452, 503, 115, 69, 497, 353]. Necessarily [167]. Necessity [113, 199, 108, 627]. Ned [225]. Need [599, 442]. Negative [44]. Neglect [605]. Neopositivismo [575]. Nersessian [428]. Neuen [520]. Neutrality [593, 594]. Newcomb [402]. Newton [123]. Newtonian [165, 181, 46]. Neyman [101, 102]. Nicholas [222, 156, 305, 410]. Nickles [152]. Nicod [65]. Niels [485, 560]. Night [464]. Niiniluoto [432]. Nineteenth [562]. Nisbett [238]. No [316, 15]. Noam [224, 89, 686]. Nominalism [125]. Nominalized [535, 340, 536]. Nomological [365]. Norma [501]. Norman [517]. Normative [45, 382, 144]. Note [361, 87, 169, 85, 188, 105, 652, 367]. Nothing [660]. Notice [275, 685, 603]. Notions [529]. Novel [189]. Nuclear [391]. Number [300]. Numbers [125]. Nussbaum [17].

Oaklander [524]. Objection [257, 701]. Objective [688]. Objectives [254]. Objectivity [692]. Objects [612, 131, 304]. Observation [292, 193, 692, 626]. Occasion [672]. Okruhlik [655]. Old [548, 332]. Olivier [705]. Olscamp [156]. One [558]. Only [332]. Ontogenetic [116]. Ontological [98, 97, 463]. Ontology [87, 196, 44]. Open [275, 518]. Operant [57, 420]. Optical [437]. Optimal [55]. Optimal-Design [55]. Options [698]. Organism [394]. Organismic [500]. Origin [242, 527, 686]. Other [405, 304]. Otherwise [474]. Otte [531]. Otto [451, 391]. Owen [626].

P [282, 627, 571]. P. [95]. Pair [249]. Palmarini [224]. Paolo [430]. Papers [258]. Papineau [122]. Paradigm [513]. Paradox [107, 702, 447, 257, 218, 364, 459, 64, 420, 582, 8, 531]. Paradoxes [706, 538]. Pargetter [67]. Parsimony [555]. Parsons [496]. Particle [427, 376, 117]. Particles [300, 706]. Particulars [197]. Partition [648]. Partition-Theorems [648]. Partly [683]. past [621, 681, 510]. Patrick [370, 639, 190]. Patterns [116]. Paty [90]. Paul [156, 448, 277, 413]. Peacocke [671]. Pearce [49, 262]. Pearson [101, 102]. Pedersen [560].

Peirce [433, 533]. Pera [279]. Percept [631]. Perception [445, 370, 48]. Perceptual [593, 594]. Peripatos [412]. Persons [148]. Perspective [15, 324, 451]. Perspectives [320]. Peter [35, 334, 626, 706, 188]. Petersen [271, 532]. Phase [668, 39]. Phenomena [195]. Phenomenological [149]. Phenomenology [530]. Phenomonon [20]. Phenotype [319]. Philip [388, 497, 317, 315]. Philosophers [123]. Philosophic [290]. Philosophical [372, 431, 585, 440, 352, 561, 156, 258]. Philosophie [244, 412]. Philosophy [155, 222, 433, 370, 259, 276, 289, 483, 291, 502, 460, 158, 398, 233, 18, 448, 409, 112, 550, 266, 389, 485, 84, 645, 203, 523, 374, 392, 519, 672, 475, 33, 225, 284, 446, 655, 193, 347, 48, 285, 441, 282, 47, 450, 496, 281, 707, 69, 413, 544]. Phylogenetic [50]. Physical [260, 705, 25, 676, 626, 318, 196, 517]. Physics [189, 406, 117, 669, 235, 275, 391, 427, 707, 408, 518, 561, 303, 259, 392, 519]. Piaget [543, 224]. Piattelli [224]. Piattelli-Palmarini [224]. Pickering [427]. Pictures [284]. Pitt [284]. Plasticity [593, 594]. Pluralism [558]. Polish [374]. Politics [559]. Popper [688, 283, 275, 443, 701, 275, 404, 683, 542, 245]. Popper-Miller [443, 683]. Popperian [455]. Population [311, 40]. Positivism [84]. Possibility [703]. Post [155]. Post-Empiricist [155]. Postulate [177, 118, 252, 324]. Poverty [455]. Power [165]. Practical [653]. Practice [77, 603]. Pragmatic [83, 451, 42, 297]. Predictability [10]. Preemption [202]. Preface [459]. Preferences [464]. Preformation [116]. Prelude [223]. Price [105]. Prices [203]. Prigogine [318]. Primer [685]. Primitive [478]. Principle [632, 188, 555, 85]. Principles [230, 414]. Priority [460]. Prism [419]. Private [302]. Probabilism [359, 638]. Probabilistic [711, 210, 439, 514, 29, 701, 614, 699, 472, 364, 639]. Probabilities [212]. Probability [58, 670, 184, 2, 41, 667, 570, 105, 400, 350, 146, 511, 507, 174, 613, 23, 277]. Problem [99, 94, 632, 676, 402, 445, 612, 154]. Problems [290, 60, 661, 393, 561, 686, 157]. Process [381]. Processes [663, 699]. Profits [203]. Program [84, 97, 455, 697]. Progress [12, 60, 474]. Progressive [432]. Projection [177, 118, 252, 324]. Prokhovnik [589]. Propensity [316, 170, 478]. Proper [665]. Properly [536]. Properties [102]. Property [534]. Proposition [614]. Provenance [483]. Provocations [240]. Proximate [423]. Pseudo [676]. Pseudo-Problem [676]. Psychiatry [393]. Psychoanalysis [440, 306, 466, 392]. Psychofunctionalism [493]. Psychogenèse [543]. Psychological [195, 503]. Psychology [232, 11, 200, 403, 386, 515, 456, 225, 382, 151, 347, 24, 144, 646, 72]. Ptolemaic [557]. Pump [516]. Punctuated [253]. Punctuationists [363, 583]. Purpose [96]. Putnam [672, 351, 413]. Pylyshyn [540].

Quality [333]. Quantitative [358, 696]. Quantum [569, 177, 182, 611, 184, 361, 85, 119, 429, 518, 118, 573, 188, 70, 275, 249, 477, 175, 548, 360, 183, 368, 697, 419, 190, 185, 267, 706, 235, 590, 275, 90, 35, 407].

Quantum-Mechanical [569]. Quarks [427, 561]. Quarrel [123]. Quarter [562]. Quasars [561]. Quentin [621]. Quest [497, 143]. Question [210, 418]. Questions [518, 235, 587]. Quests [473]. Quine [285, 450].

R [689, 283, 15, 275, 559, 392]. Rachel [429]. Radiation [548]. Radical [257, 372]. Rainer [50]. Random [272]. Randomization [571, 384]. Rare [534]. Rational [237, 595, 150, 23, 468]. Rationality [618, 585, 596, 574, 438, 475, 584, 418, 619, 294, 231, 654]. Readings [189, 225]. Real [600, 301, 280]. Realism [169, 567, 198, 78, 301, 438, 4, 530, 267, 590, 77, 405, 283]. Realist [409, 230]. Reality [695, 265, 221]. Reason [143, 77, 34, 387, 351]. Reasoning [447, 59, 24, 299, 521]. Reasons [595, 201]. Received [178, 246, 321, 395, 469, 545, 606, 673]. Reconceptualizations [312]. Reconsidered [292, 403]. Reconstructions [276]. Red [137]. Redhead [556]. Reduce [11, 151]. Reduction [98, 491, 200, 149, 342, 652, 473, 651, 221]. Redundancy [104]. Reference [601, 61, 250, 54, 143]. Referral [94, 95]. Regress [445]. Reichenbach [153, 425]. Reinstatement [556]. Reinterpretation [501]. Rejecting [539]. Rejoinder [711]. Relation [472, 482]. Relational [229]. Relationalism [602]. Relationist [696]. Relations [524, 688, 211, 456, 472]. Relative [600]. Relativity [622, 425, 359, 638, 476, 589, 261]. Relevance [255, 94, 554]. Reliability [596]. Remark [425]. Remote [681]. Renormalization [662]. Rensch [500]. Replies [222]. Reply [98, 120, 593, 512, 599, 514, 617, 594, 570, 344, 10, 95, 477, 86, 172, 274, 366]. Representation [41, 341, 651]. Representational [680]. Representations [89]. Representing [389]. Require [613]. Requirements [585]. Requires [12]. Rescher [222, 8, 305, 410]. Research [15, 403, 84, 473]. Resistance [26]. Resolution [457]. Response [556, 330, 171, 534, 619]. Retrievable [698]. Review [155, 222, 90, 260, 524, 405, 336, 705, 656, 244, 486, 433, 688, 370, 375, 391, 205, 35, 189, 259, 468, 276, 604, 623, 157, 372, 407, 483, 541, 280, 689, 588, 502, 20, 431, 427, 71, 640, 690, 707, 304, 692, 242, 282, 487, 440, 671, 221, 428, 158, 139, 108, 334, 572, 501, 480, 406, 408, 624, 125, 258, 451, 687, 390, 625, 574, 18, 49, 394, 376, 448, 704, 605, 47, 173, 409, 32, 540, 283, 429, 518, 91, 333, 657, 371, 15, 626, 318, 542, 573, 220, 302, 352, 19, 153, 36, 414, 305]. Review [335, 589, 124, 543, 520, 450, 430, 306, 393, 466, 123, 17, 389, 432, 410, 639, 69, 176, 152, 245, 521, 627, 154, 485, 388, 350, 224, 515, 241, 240, 500, 70, 561, 203, 16, 223, 73, 377, 559, 523, 121, 277, 72, 261, 374, 392, 519, 672, 562, 175, 122, 204, 387, 262, 351, 243, 89, 522, 50, 33, 225, 319, 354, 284, 497, 34, 138, 317, 498, 320, 655, 413, 279, 411, 412, 575, 278, 68, 137, 353, 174, 449, 202, 369, 190, 706, 303, 641, 654, 481, 337, 496, 590, 691, 48, 560, 373, 452]. Review [503, 499, 465, 516, 467, 517, 285, 484, 482, 544]. Reviews [281, 156, 686]. Revisions [599, 270, 442]. Revisited [119, 582]. Revolution [137, 598]. Revolutions [276]. Reward [19]. Rhetoric [559]. Richard [604, 221, 69, 319, 138, 174]. Richards [690]. Rid [218]. Riddle [410, 134]. Right [201]. Rigid [537]. Rise [391]. Robert [336, 433, 189, 690, 480, 49, 241, 561, 377, 262, 411, 68, 484, 346]. Roger [285]. Rolando [543]. Role [45, 589, 39, 481]. Romano [522]. Romanos [450]. Room [691]. Rorty [139, 69]. Rosenberg [487, 330, 446, 274, 202]. Rossi [430]. Rotation [425]. Route [568]. Rowan [243]. Rudwick [499]. Rule [79, 111, 185]. Rules [238, 302, 627, 385, 89]. Rupert [123]. Ruse [137]. Russell [44, 426, 669]. Russellian [524]. Ruth [405].

S [336, 433, 205, 158, 376, 589, 95, 627, 241, 377, 392, 175, 33, 411, 499, 482]. Salmon [153, 572]. Sandra [91]. Sarkar [390]. Saul [302]. Scepticism [467]. Schaefer [652]. Schaffer [516]. Schism [275]. Schlick [258, 258]. Schoeman [512]. Schroeder [652]. Schroeder-Heister [652]. Schuster [559]. Schwartz [353]. Science [5, 155, 166, 618, 289, 483, 291, 460, 3, 295, 687, 398, 49, 535, 228, 550, 340, 536, 194, 432, 479, 515, 645, 97, 551, 475, 33, 325, 354, 273, 256, 343, 317, 697, 193, 418, 137, 353, 481, 155, 370, 276, 372, 502, 20, 71, 487, 409, 540, 283, 626, 19, 389, 241, 73, 523, 522, 284, 68, 202, 48, 484, 375, 125, 241, 204, 481, 467]. Sciences [543, 625, 318, 542, 393, 377, 374, 411]. Scientific [552, 336, 585, 428, 501, 173, 414, 438, 521, 186, 254, 16, 559, 328, 530, 678, 481, 499, 77, 603, 572, 408, 152, 278]. Scientist [36]. Scope [114]. Scoring [385]. Scott [529]. Search [156, 387]. Searle [462]. Seit [520]. Selected [259, 240]. Selection [607, 309, 83, 431, 650, 596, 214, 677, 566, 549, 616, 239, 682, 161, 319]. Selectionism [274]. Sellar [284]. Sellarsian [164]. Selleri [376]. Semantic [567, 311, 328]. Semantics [345, 612, 343]. Sensations [120]. Sensory [95]. Sergio [688]. Set [623, 335]. Sets [2, 315]. Seventy [519]. Seventy-Fifth [519]. Shaky [590]. Shall [162]. Shames [64]. Shapere [387]. Shapin [516]. Shaping [499]. Shapiro [657]. Shea [391]. Shimony [519]. Siegel [618, 256]. Significance [483]. Simon [516, 163]. Simpson [531, 364]. Simultaneity [67]. Sir [243]. Size [623]. Sklar [707]. Skyrms [108]. Slippery [680]. Slope [680]. Smith [621]. Snizek [15]. Sober [431, 632, 344, 534, 616]. Social [155, 377, 202, 77]. Society [241]. Sociobiology [497, 513, 202]. Sociological [427]. Sociology [15, 241, 414]. Solipsism [403, 386, 646]. Solution [134]. Solve [670, 444]. Solving [661]. Some [703, 60, 249, 342, 235, 583]. Sorabji [604]. Sosa [222]. Sources [522]. Space [602, 196, 261, 229, 437, 33, 187, 465, 46, 370, 242]. Space-Perception [370]. Spacetime [707, 229]. Spalek [91]. Speaking [91]. Special [261, 622, 359, 638, 476]. Specialists [499]. Species [87, 617, 313, 558, 88, 314, 527, 315, 610, 421]. Specificity [107]. Speed [703]. Spinoza [625]. Split [148]. Squire [689]. Stable [698]. Stachel [33]. Stalnaker [262, 480]. Standards [512]. Standing [306]. Starr [320]. Statements [2]. States [91]. Statistical [66, 565, 417, 80]. Statistics [401, 419]. Status [380]. Stegmüller [672, 520]. Steiner [669]. Stephen [515, 641]. Steven [516]. Stewart [657]. Stich [238, 515]. Still [637]. Stillman [16]. Stochastic [568]. Stove [574, 245]. Strange [56, 249]. Strategy [55, 403, 527]. Structural [677]. Structure [281, 487, 572, 25, 311, 249, 130, 527, 343, 681]. Structures [649]. Strukturalismus [520]. Stuart [112]. Studies [205, 241, 33, 152, 354, 174, 190]. Study [232, 303]. Subjective [168, 95]. Subjunctive [367]. Substitute [316]. Substitution [183]. Substitutional [345]. Substitutivity [105]. Substrata [197]. Subsumption [213]. Subtler [65]. Suggestions [528]. Sum [79]. Suppes [639, 190]. Support [556, 614, 214, 683]. Surrounding [342]. Swartz [517]. Symbiosis [607, 549]. Symmetry [611]. Synonymy [187]. Syntactic [700]. Syntax [612]. Synthesis [472, 363, 583]. Synthetic [253]. System [19]. Systematics [598, 50]. Systems [345]. Szabo [281].

T [656, 352, 175, 320]. Tarozzi [518]. Taxonomy [362]. Tears [663]. Teasers [670]. Technology [5]. Teleological [217]. Teleology [329, 57, 420]. Tempo [253]. Temporal [524, 524]. Tendencies [234]. Tense [367]. Terms [54]. Test [404]. Testability [1, 171]. Testing [599, 270, 417, 442, 603]. Tests [79]. Text [17]. Their [481]. Themes [371]. Theorem [667]. Theorems [648]. Theoretic [4, 508]. Theoretical [7, 593, 594, 327, 54, 421]. Theories [205, 604, 650, 690, 631, 257, 580, 254, 196, 27, 229, 696, 33, 463, 268, 651, 648, 300, 428, 465, 603]. Theory [557, 552, 524, 688, 623, 361, 407, 431, 458, 700, 29, 327, 390, 119, 15, 402, 335, 266, 579, 194, 105, 214, 26, 39, 101, 275, 424, 562, 175, 548, 104, 349, 43, 423, 172, 587, 253, 48, 476, 115, 484, 486, 371, 590, 73, 121, 122, 337]. There [58, 211]. Thermodynamics [80, 468]. Thermostatics [149]. Thesis [331]. Think [65]. Thinking [40]. Third [296]. Thomas [156, 152, 243, 320, 155, 71]. Thorndike [136, 1]. Thought [405, 462]. Thoughts [671]. Tile [538]. Time [611, 622, 221, 318, 218, 430, 95, 602, 196, 261, 437, 33, 187, 465, 46, 524, 280, 262, 705, 604, 352]. Timing [120, 168]. Tisza [519]. Tolman [232]. Tooth [137]. Topics [389, 560]. Toraldo [260]. Total [667]. Tradition [560]. Traditional [479]. Transcendental [451]. Transcendental-Pragmatic [451]. Transformation [476]. Transitivity [210]. Translation [17, 652]. Trial [102]. Tribute [376]. Trivialization [66]. Trouble [27]. Truer [634]. Truesdell [468]. Truth [704, 660, 186, 351, 183, 133, 156, 156]. Truthlikeness [455]. Turchin [20]. Turing [63]. Twentieth [371]. Twentieth-Century [371]. Two [195, 614, 537, 633]. Two-Proposition [614]. Tycho [483]. Type [650, 171]. Types [62, 273].

Ultimate [423]. Unanswered [235]. Uncaused [581]. Uncertainty [361, 188, 490, 85]. Underdetermination [328]. Understanding [484, 451, 261]. Ungar [281]. Unification [98, 128]. Unified [349, 463, 300]. Uniqueness [602, 28]. Unit [319]. United [91]. Units [607, 83, 677, 549].

Unity [97, 697, 68]. Universe [215, 589, 681, 332, 581, 275]. Unnoticed [134]. Unrevised [441]. Unscientific [630]. Urbach [571]. Ursus [483]. Use [686]. Using [529]. Utility [633].

V [285]. Vague [2]. Valentin [20]. Valid [299]. Value [368, 267]. Values [6, 325, 481]. Vance [242]. Variable [232]. Variables [211, 568]. Variance [257]. Variant [404]. Varieties [691]. Vaulting [497]. Vector [631]. Velde [258]. Velde-Schlick [258]. Verisimilitude [162]. versus [669, 461]. Vico [430]. Vision [579]. Vitamins [312]. Vol [225]. Volume [75, 141, 207, 287, 356, 435, 505, 577, 643, 709, 70, 33, 174].

W [391, 540, 283, 573, 241, 377, 34, 411, 285]. Wallace [429, 414, 522, 223]. Walter [414]. Wanting [691]. War [123]. Warnke [451]. Warren [18]. Wartofsky [241, 377, 411]. Watkins [467]. Wave [376]. Wave-Particle [376]. Weber [502]. Wesley [572, 153]. Wheeler [407]. Whewell [383]. Who [641]. Why-Questions [587]. Wide [362]. Wilfrid [284]. Wilhelm [672]. Will [691]. Willi [50]. William [157, 640, 15, 223, 262, 243, 522, 641]. Williams [617]. Wilson [541]. Wissenschaftstheorie [244]. within [181]. without [125, 647, 301, 663, 476]. Witkowski [656]. Wittgenstein [302, 124, 627]. Wladyslaw [374]. Wojciech [407]. Wolfgang [520, 672]. Woodfield [329]. Work [394, 39, 16]. World [260, 572]. worth [691]. Would [366]. Wright [124, 217]. Writings [240]. Wrong [700, 363, 583]. Wylie [656].

Years [291]. Yeo [559].

Z [562]. Zabell [668]. Zangerl [50]. Zeno [582, 538]. Zenon [540]. Zurek [407].

References

Laymon:1980:ITM

[1] Ronald Laymon. Independent testability: The Michelson–Morley and Kennedy–Thorndike experiments. Philosophy of Science, 47(1):1–37, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

     tronic). URL http://www.jstor.org/stable/187140.

Dale:1980:PVS

[2] A. I. Dale. Probability, vague statements and fuzzy sets. Philosophy of Science, 47(1):38–55, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

     187141.

Elgin:1980:LES

[3] C. Z. Elgin. Lawlikeness and the end of science. Philosophy of Science, 47(1):56–68, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187142.

Merrill:1980:MTA

[4] G. H. Merrill. The model-theoretic argument against realism. Philosophy of Science, 47(1):69–81, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187143.

Agassi:1980:BST

[5] Joseph Agassi. Between science and technology. Philosophy of Science, 47(1):82–99, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187144.

Goosens:1980:VHM

[6] William K. Goosens. Values, health, and medicine. Philosophy of Science, 47(1):100–115, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187145.

Bunzl:1980:CHT

[7] Martin Bunzl. Comment on “health as a theoretical concept”. Philosophy of Science, 47(1):116–118, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187146.

Bunch:1980:RGP

[8] B. L. Bunch. Rescher on the Goodman Paradox. Philosophy of Science, 47(1):119–123, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187147.

Dieks:1980:ECD

[9] D. Dieks. On the empirical content of determinism. Philosophy of Science, 47(1):124–130, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187148.

Kukla:1980:DPR

[10] A. Kukla. Determinism and predictability: Reply to Dieks. Philosophy of Science, 47(1):131–133, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187149.

Kitcher:1980:HRF

[11] Patricia Kitcher. How to reduce a functional psychology? Philosophy of Science, 47(1):134–140, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187150.

Kordig:1980:PRI

[12] Carl R. Kordig. Progress requires invariance. Philosophy of Science, 47(1): 141, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187151.

Garber:1980:FJC

[13] Daniel Garber. Field and Jeffrey conditionalization. Philosophy of Science, 47(1):142–145, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187152.

Adler:1980:WMB

[14] Carl G. Adler. Why is mechanics based on acceleration? Philosophy of Science, 47(1):146–152, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187153.

Homans:1980:BRB

[15] George C. Homans. Book review: Contemporary Issues in Theory and Research: A Metasociological Perspective“ (Contributions in Sociology, No.

33) by William E. Snizek; Ellsworth R. Fuhrman; Michael K. Miller. Philosophy of Science, 47(1):153–154, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187154.

McMullin:1980:BRB

[16] Ernan McMullin. Book review: Galileo at Work: His Scientific Biography by Stillman Drake. Philosophy of Science, 47(1):154–156, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187155.

Lennox:1980:BRB

[17] James G. Lennox. Book review: Aristotle’s De Motu Animalium: Text with Translation, Commentary and Interpretive Essays by Martha Craven Nussbaum. Philosophy of Science, 47(1):156–159, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187156.

Graham:1980:BRB

[18] Colin C. Graham. Book review: Georg Cantor, His Mathematics and Philosophy of the Infinite by Joseph Warren Dauben. Philosophy of Science, 47(1):159–160, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187157.

Hull:1980:BRB

[19] David L. Hull. Book review: The Reward System in British and American Science by Jerry Gaston. Philosophy of Science, 47(1):160–161, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187158.

Colodny:1980:BRB

[20] Robert G. Colodny. Book review: The Phenomonon of Science by Valentin F. Turchin; B. Frantz. Philosophy of Science, 47(1):162, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187159.

Anonymous:1980:BMa

[21] Anonymous. Back matter. Philosophy of Science, 47(1):163–164, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187160.

Anonymous:1980:FMa

[22] Anonymous. Front matter. Philosophy of Science, 47(1):??, March 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187139.

VanFraassen:1980:RBP

[23] Bas C. Van Fraassen. Rational belief and probability kinematics. Philosophy of Science, 47(2):165–187, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187082.

Stich:1980:JPH

[24] Stephen P. Stich and Richard E. Nisbett. Justification and the psychology of human reasoning. Philosophy of Science, 47(2):188–202, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187083.

Forge:1980:SPE

[25] John Forge. The structure of physical explanation. Philosophy of Science, 47(2):203–226, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187084.

Lugg:1980:TCR

[26] Andrew Lugg. Theory choice and resistance to change. Philosophy of Science, 47(2):227–243, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187085.

Margolis:1980:THT

[27] Joseph Margolis. The Trouble with homunculus theories. Philosophy of Science, 47(2):244–259, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187086.

Fales:1980:UHL

[28] Evan Fales. Uniqueness and historical laws. Philosophy of Science, 47(2): 260–276, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187087.

Falmagne:1980:PTE

[29] Jean-Claude Falmagne. A probabilistic theory of extensive measurement. Philosophy of Science, 47(2):277–296, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187088.

Siegel:1980:JDN

[30] Harvey Siegel. Justification, discovery and the naturalizing of epistemology. Philosophy of Science, 47(2):297–321, June 1980. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187089.

Glymour:1980:HDH

[31] Clark Glymour. Hypothetico-deductivism is hopeless. Philosophy of Science, 47(2):322–325, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187090. See corrections [51].

Haugeland:1980:BRB

[32] John Haugeland. Book review: Brainstorms by Daniel C. Dennett. Philosophy of Science, 47(2):326–327, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187091.

Quinn:1980:BRB

[33] Philip L. Quinn. Book review: Foundations of Space–time Theories (Minnesota Studies in the Philosophy of Science, Volume 8) by John S. Earman; Clark N. Glymour; John J. Stachel. Philosophy of Science, 47(2): 327–329, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187092.

Rosenkrantz:1980:BRB

[34] Roger D. Rosenkrantz. Book review: Chance, Cause, Reason by Arthur W. Burks. Philosophy of Science, 47(2):329–332, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187093.

Bub:1980:BRB

[35] Jeffrey Bub. Book review: Quantum Logic by Peter Mittelstaedt. Philosophy of Science, 47(2):332–335, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187094.

Ihde:1980:BRB

[36] Aaron J. Ihde. Book review: Gay-Lussac: Scientist and Bourgeois by Maurice Crosland. Philosophy of Science, 47(2):335–336, June 1980. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187095.

Anonymous:1980:BMb

[37] Anonymous. Back matter. Philosophy of Science, 47(2):337–338, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187096.

Anonymous:1980:FMb

[38] Anonymous. Front matter. Philosophy of Science, 47(2):??, June 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187081.

Malament:1980:WGP

[39] David B. Malament and Sandy L. Zabell. Why Gibbs phase averages work — the role of ergodic theory. Philosophy of Science, 47(3):339–349, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186949. See comment

[668].

Sober:1980:EPT

[40] Elliott Sober. Evolution, population thinking, and essentialism. Philosophy of Science, 47(3):350–383, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186950.

Domotor:1980:PKR

[41] Zoltan Domotor. Probability kinematics and representation of belief change. Philosophy of Science, 47(3):384–403, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186951.

Gardenfors:1980:PAE

[42] Peter Gardenfors. A pragmatic approach to explanations. Philosophy of Science, 47(3):404–423, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186952.

Roberts:1980:LTM

[43] Fred S. Roberts. On Luce’s theory of meaningfulness. Philosophy of Science, 47(3):424–433, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186953.

Oaklander:1980:RNF

[44] L. Nathan Oaklander and Silvano Miracchi. Russell, negative facts, and ontology. Philosophy of Science, 47(3):434–455, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186954.

Currie:1980:RNA

[45] Gregory Currie. The role of normative assumptions in historical explanation. Philosophy of Science, 47(3):456–473, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186955.

Zaret:1980:LCC

[46] David Zaret. A limited conventionalist critique of Newtonian space–time. Philosophy of Science, 47(3):474–494, September 1980. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186956.

Hankins:1980:BRB

[47] Thomas L. Hankins. Book review: The Concept of Matter in Modern Philosophy by Ernan McMullin. Philosophy of Science, 47(3):495–496, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186957.

Townsend:1980:BRB

[48] Burke Townsend. Book review: Perception, Theory and Commitment. The New Philosophy of Science by Harold I. Brown. Philosophy of Science, 47 (3):496–498, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186958.

Grant:1980:BRB

[49] Edward Grant. Book review: From Humanism to Science 1480–1700 by Robert Mandrou; Brian Pearce. Philosophy of Science, 47(3):498–499, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186959.

Platnick:1980:BRB

[50] Norman I. Platnick and Gareth Nelson. Book review: Phylogenetic Systematics by Willi Hennig; D. Dwight Davis; Rainer Zangerl. Philosophy of Science, 47(3):499–502, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186960.

Anonymous:1980:CHD

[51] Anonymous. Corrections: Hypothetico-deductivism is hopeless. Philosophy of Science, 47(3):503, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186961. See [31].

Anonymous:1980:BMc

[52] Anonymous. Back matter. Philosophy of Science, 47(3):504, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186962.

Anonymous:1980:FMc

[53] Anonymous. Front matter. Philosophy of Science, 47(3):??, September 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186948.

Nola:1980:FRT

[54] Robert Nola. Fixing the reference of theoretical terms. Philosophy of Science, 47(4):505–531, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187311.

Beatty:1980:ODM

[55] John Beatty. Optimal-design models and the strategy of model building in evolutionary biology. Philosophy of Science, 47(4):532–561, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187312.

Hacking:1980:SE

[56] Ian Hacking. Strange expectations. Philosophy of Science, 47(4):562–567, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187313.

Porpora:1980:OCT

[57] Douglas V. Porpora. Operant conditioning and teleology. Philosophy of Science, 47(4):568–582, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187314.

Armendt:1980:TDB

[58] Brad Armendt. Is there a Dutch book argument for probability kinematics? Philosophy of Science, 47(4):583–588, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187315.

Fumerton:1980:IRB

[59] R. A. Fumerton. Induction and reasoning to the best explanation. Philosophy of Science, 47(4):589–600, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187316.

Krips:1980:SPP

[60] H. Krips. Some problems for “progress and its problems”. Philosophy of Science, 47(4):601–616, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187317.

Hesse:1980:MIR

[61] Mary Hesse. Minogue on intensional reference. Philosophy of Science, 47 (4):617–625, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187318.

Cupples:1980:FTE

[62] Brian Cupples. Four types of explanation. Philosophy of Science, 47(4): 626–629, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187319.

Nelson:1980:TMA

[63] R. J. Nelson. Turing machine arguments. Philosophy of Science, 47(4): 630–633, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187320.

Primeaux:1980:SEE

[64] D. Primeaux. On Shames’ experimenter expectancy paradox. Philosophy of Science, 47(4):634–637, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187321.

Rozeboom:1980:NCS

[65] William W. Rozeboom. Nicod’s criterion: Subtler than you think. Philosophy of Science, 47(4):638–643, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187322.

Cooke:1980:TND

[66] Roger Cooke. A trivialization of Nagel’s definition of explanation for statistical laws. Philosophy of Science, 47(4):644–645, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187323.

Townsend:1980:JPD

[67] Burke Townsend. Jackson and Pargetter on distant simultaneity. Philosophy of Science, 47(4):646–655, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187324.

Sklar:1980:BRB

[68] Lawrence Sklar. Book review: Unity of Science by Robert L. Causey. Philosophy of Science, 47(4):656–657, December 1980. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187325.

Lieberson:1980:BRB

[69] Jonathan Lieberson. Book review: Philosophy and the Mirror of Nature by Richard Rorty. Philosophy of Science, 47(4):657–659, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187326.

McGrath:1980:BRB

[70] James H. McGrath. Book review: The Logico-Algebraic Approach to Quantum Mechanics. Volume II: Contemporary Consolidation by C. A. Hooker. Philosophy of Science, 47(4):660–662, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187327.

Darden:1980:BRB

[71] Lindley Darden. Book review: Thomas Hunt Morgan, The Man and His Science by Garland E. Allen. Philosophy of Science, 47(4):662–666, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187328.

Moskowitz:1980:BRB

[72] Merle J. Moskowitz. Book review: The First Century of Experimental Psychology by Eliot Hearst. Philosophy of Science, 47(4):666–667, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187329.

Merrill:1980:BRB

[73] G. H. Merrill. Book review: Theory of Science by George Gale. Philosophy of Science, 47(4):667–670, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187330.

Anonymous:1980:BMd

[74] Anonymous. Back matter. Philosophy of Science, 47(4):671–679, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187331.

Anonymous:1980:VI

[75] Anonymous. Volume information. Philosophy of Science, 47(4):673–678, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187309.

Anonymous:1980:FMd

[76] Anonymous. Front matter. Philosophy of Science, 47(4):??, December 1980. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187310.

Will:1981:RSP

[77] Frederick L. Will. Reason, social practice, and scientific realism. Philosophy of Science, 48(1):1–18, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187065.

Laudan:1981:CCR

[78] Larry Laudan. A confutation of convergent realism. Philosophy of Science, 48(1):19–49, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187066.

Redhead:1981:ETS

[79] M. L. G. Redhead. Experimental tests of the sum rule. Philosophy of Science, 48(1):50–64, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187067.

Wicken:1981:CEC

[80] Jeffrey S. Wicken. Causal explanations in classical and statistical thermodynamics. Philosophy of Science, 48(1):65–77, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187068.

Leeds:1981:KFI

[81] Stephen Leeds. Kyburg and fiducial inference. Philosophy of Science, 48(1):78–91, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187069.

Kyburg:1981:LIM

[82] Henry E. Kyburg, Jr. Leeds’ infernal machine. Philosophy of Science, 48(1):92–94, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187070.

Cassidy:1981:APF

[83] John Cassidy. Ambiguities and pragmatic factors in the units of selection controversy. Philosophy of Science, 48(1):95–111, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187071.

Martin:1981:BPR

[84] Raymond Martin. Beyond positivism: A research program for philosophy of history. Philosophy of Science, 48(1):112–121, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187072.

Gibbins:1981:NQL

[85] Peter Gibbins. A note on quantum logic and the Uncertainty Principle. Philosophy of Science, 48(1):122–126, March 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187073. See note [188].

Nissim-Sabat:1981:RGJ

[86] Charles Nissim-Sabat. A reply to Grunbaum and Janis. Philosophy of Science, 48(1):127–129, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187074.

Caplan:1981:BCN

[87] Arthur L. Caplan. Back to class: A note on the ontology of species. Philosophy of Science, 48(1):130–140, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187075.

Hull:1981:KKC

[88] David L. Hull. Kitts and Kitts and Caplan on species. Philosophy of Science, 48(1):141–152, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187076.

Perfetti:1981:BRB

[89] Charles A. Perfetti. Book review: Rules and Representations by Noam Chomsky. Philosophy of Science, 48(1):153–156, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187077.

Arthur:1981:BRB

[90] Richard T. W. Arthur. Book review: Quantum Mechanics, a Half Century Later by J.L. Lopes; M. Paty. Philosophy of Science, 48(1):156–161, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187078.

Heilbron:1981:BRB

[91] J. L. Heilbron. Book review: Guide to the Archival Materials of the German-Speaking Emigration to the United States after 1933 by John M. Spalek; Adrienne Ash; Sandra H. Hawrylchak. Philosophy of Science, 48 (1):161, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187079.

Anonymous:1981:BMa

[92] Anonymous. Back matter. Philosophy of Science, 48(1):162–164, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187080.

Anonymous:1981:FMa

[93] Anonymous. Front matter. Philosophy of Science, 48(1):??, March 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187064.

Churchland:1981:ABR

[94] Patricia Smith Churchland. On the alleged backwards referral of experiences and its relevance to the mind-body problem. Philosophy of Science, 48(2):165–181, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187178.

See reply [95].

Libet:1981:EES

[95] Benjamin Libet. The experimental evidence for subjective referral of a sensory experience backwards in time: Reply to P. S. Churchland. Philosophy of Science, 48(2):182–197, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187179. See [94].

Falk:1981:PFE

[96] Arthur E. Falk. Purpose, feedback, and evolution. Philosophy of Science, 48(2):198–217, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187180.

McCauley:1981:HIO

[97] Robert N. McCauley. Hypothetical identities and ontological economizing: Comments on Causey’s program for the unity of science. Philosophy of Science, 48(2):218–227, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187181.

Causey:1981:ROU

[98] Robert L. Causey. Reduction and ontological unification: Reply to Mc-Cauley. Philosophy of Science, 48(2):228–231, June 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187182.

Christensen:1981:PI

[99] Ferrel Christensen. The problem of inertia. Philosophy of Science, 48(2): 232–247, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187183.

Cherniak:1981:FI

[100] Christopher Cherniak. Feasible inferences. Philosophy of Science, 48(2): 248–268, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187184.

Mayo:1981:DNP

[101] Deborah G. Mayo. In defense of the Neyman–Pearson theory of confidence intervals. Philosophy of Science, 48(2):269–280, June 1981. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187185.

Seidenfeld:1981:ATP

[102] Teddy Seidenfeld. On after-trial properties of best Neyman–Pearson confidence intervals. Philosophy of Science, 48(2):281–291, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187186.

Edidin:1981:GC

[103] Aron Edidin. Glymour on confirmation. Philosophy of Science, 48(2):292– 307, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187187.

Partridge:1981:ITR

[104] Derek Partridge. Information theory and redundancy. Philosophy of Science, 48(2):308–316, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187188.

Leblanc:1981:WPS

[105] Hugues Leblanc. What price substitutivity? A note on probability theory. Philosophy of Science, 48(2):317–322, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187189.

Lennox:1981:EHC

[106] James Lennox. Enc on Harvey and consequence etiologies. Philosophy of Science, 48(2):323–326, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187190.

Cooke:1981:PHC

[107] Roger M. Cooke. A paradox in Hempel’s criterion of maximal specificity. Philosophy of Science, 48(2):327–328, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187191.

Fetzer:1981:BRB

[108] James H. Fetzer. Book review: Causal Necessity by Brian Skyrms. Philosophy of Science, 48(2):329–335, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187192.

Anonymous:1981:BMb

[109] Anonymous. Back matter. Philosophy of Science, 48(2):??, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187193.

Anonymous:1981:FMb

[110] Anonymous. Front matter. Philosophy of Science, 48(2):??, June 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187177.

Shafer:1981:JRC

[111] Glenn Shafer. Jeffrey’s Rule of Conditioning. Philosophy of Science, 48 (3):337–362, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186984.

Hausman:1981:JSM

[112] Daniel M. Hausman. John Stuart Mill’s philosophy of economics. Philosophy of Science, 48(3):363–385, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/186985. See comment [441].

Hochberg:1981:NNL

[113] Herbert Hochberg. Natural necessity and laws of nature. Philosophy of Science, 48(3):386–399, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186986.

Kitchener:1981:NSG

[114] Richard F. Kitchener. The nature and Scope of genetic epistemology. Philosophy of Science, 48(3):400–415, September 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186987.

Wassermann:1981:NTE

[115] Gerhard D. Wassermann. On the nature of the theory of evolution. Philosophy of Science, 48(3):416–437, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186988.

Katz:1981:POP

[116] Michael J. Katz and William Goffman. Preformation of ontogenetic patterns. Philosophy of Science, 48(3):438–453, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186989.

Hendrick:1981:AIC

[117] R. E. Hendrick and Anthony Murphy. Atomism and the illusion of crisis: The danger of applying Kuhnian categories to current particle physics. Philosophy of Science, 48(3):454–468, September 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186990.

Hellman:1981:QLP

[118] Geoffrey Hellman. Quantum logic and the projection postulate. Philosophy of Science, 48(3):469–486, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/186991. See corrections [177].

Ginsberg:1981:QTI

[119] Allen Ginsberg. Quantum theory and the identity of indiscernibles revisited. Philosophy of Science, 48(3):487–491, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186992.

Churchland:1981:TSR

[120] Patricia Smith Churchland. The timing of sensations: Reply to Libet. Philosophy of Science, 48(3):492–497, September 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186993.

Morton:1981:BRB

[121] Adam Morton. Book review: Theory and Evidence by Clark Glymour. Philosophy of Science, 48(3):498–500, September 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186994.

Nickles:1981:BRB

[122] Thomas Nickles. Book review: Theory and Meaning by David Papineau. Philosophy of Science, 48(3):500–502, September 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186995.

Lennon:1981:BRB

[123] Thomas M. Lennon. Book review: Philosophers at War: The Quarrel between Newton and Leibniz by A. Rupert Hall. Philosophy of Science, 48 (3):502–503, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186996.

Kielkopf:1981:BRB

[124] Charles F. Kielkopf. Book review: Wittgenstein on the Foundations of Mathematics by Crispin Wright. Philosophy of Science, 48(3):503–505, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186997.

Friedman:1981:BRB

[125] Michael Friedman. Book review: Science without Numbers: A Defense of Nominalism by Hartry H. Field. Philosophy of Science, 48(3):505–506, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186998.

Anonymous:1981:BMc

[126] Anonymous. Back matter. Philosophy of Science, 48(3):??, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186999.

Anonymous:1981:FMc

[127] Anonymous. Front matter. Philosophy of Science, 48(3):??, September 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186983.

Kitcher:1981:EU

[128] Philip Kitcher. Explanatory unification. Philosophy of Science, 48(4): 507–531, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186834.

Levi:1981:DIC

[129] Isaac Levi. Direct inference and confirmational conditionalization. Philosophy of Science, 48(4):532–552, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186835.

Muscari:1981:SMD

[130] Paul G. Muscari. The structure of mental disorder. Philosophy of Science, 48(4):553–572, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186836.

Sacksteder:1981:HGO

[131] William Sacksteder. Hobbes: Geometrical objects. Philosophy of Science, 48(4):573–590, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186837.

Hesslow:1981:CD

[132] Germund Hesslow. Causality and determinism. Philosophy of Science, 48 (4):591–605, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186838.

Schwartz:1981:ATC

[133] Robert John Schwartz. Approximate truth and confirmation. Philosophy of Science, 48(4):606–610, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186839.

Shirley:1981:UFB

[134] Edward S. Shirley. An unnoticed flaw in Barker and Achinstein’s solution to Goodman’s new riddle of induction. Philosophy of Science, 48(4):611– 617, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186840.

Laudan:1981:AA

[135] Larry Laudan. Anomalous anomalies. Philosophy of Science, 48(4):618– 619, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186841.

Erlichson:1981:MMK

[136] Herman Erlichson. The Michelson–Morley and Kennedy-Thorndike experiments. Philosophy of Science, 48(4):620–622, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186842. See comments [171].

Sloan:1981:BRB

[137] Phillip R. Sloan. Book review: The Darwinian Revolution: Science Red in Tooth and Claw by Michael Ruse. Philosophy of Science, 48(4):623–627, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186843.

Ruse:1981:BRB

[138] Michael Ruse. Book review: Darwinism and Human Affairs by Richard D. Alexander. Philosophy of Science, 48(4):627–628, December 1981. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186844.

Farrell:1981:BRB

[139] Daniel M. Farrell. Book review: Explaining Emotions by Amelie Rorty. Philosophy of Science, 48(4):629–630, December 1981. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186845.

Anonymous:1981:BMd

[140] Anonymous. Back matter. Philosophy of Science, 48(4):631–637, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186846.

Anonymous:1981:VI

[141] Anonymous. Volume information. Philosophy of Science, 48(4):633–635, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186832.

Anonymous:1981:FMd

[142] Anonymous. Front matter. Philosophy of Science, 48(4):??, December 1981. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186833.

Shapere:1982:RRQ

[143] Dudley Shapere. Reason, reference, and the quest for knowledge. Philosophy of Science, 49(1):1–23, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186877.

Thagard:1982:DNP

[144] Paul Thagard. From the descriptive to the normative in psychology and logic. Philosophy of Science, 49(1):24–42, March 1982. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186878.

Skyrms:1982:CDL

[145] Brian Skyrms. Counterfactual definiteness and local causation. Philosophy of Science, 49(1):43–50, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186879.

Plato:1982:PD

[146] Jan Von Plato. Probability and determinism. Philosophy of Science, 49 (1):51–66, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186880.

Fales:1982:NKF

[147] Evan Fales. Natural kinds and freaks of nature. Philosophy of Science, 49(1):67–90, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186881.

Moor:1982:SBA

[148] James Moor. Split brains and atomic persons. Philosophy of Science, 49(1):91–106, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186882.

Kuipers:1982:RPK

[149] Theo A. F. Kuipers. The reduction of phenomenological to kinetic Thermostatics. Philosophy of Science, 49(1):107–119, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186883.

Schagrin:1982:FR

[150] Morton L. Schagrin. The failure to be rational. Philosophy of Science, 49(1):120–124, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186884.

Richardson:1982:HRF

[151] Robert C. Richardson. How not to reduce a functional psychology. Philosophy of Science, 49(1):125–137, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186885.

Lugg:1982:BRB

[152] Andrew Lugg. Book review: Scientific Discovery: Case Studies by Thomas Nickles. Philosophy of Science, 49(1):138–140, March 1982. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186886.

Humphreys:1982:BRB

[153] Paul Humphreys. Book review: Hans Reichenbach: Logical Empiricist by Wesley Salmon. Philosophy of Science, 49(1):140–142, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186887.

Lycan:1982:BRB

[154] William G. Lycan. Book review: Metaphysics and the Mind-Body Problem by Michael E. Levin. Philosophy of Science, 49(1):142–144, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186888.

Airaksinen:1982:BRB

[155] Timo Airaksinen. Book review: Naturalism and Social Science: A Post-Empiricist Philosophy of Social Science by David Thomas. Philosophy of Science, 49(1):144–146, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186889.

Garber:1982:BRB

[156] Daniel Garber. Book reviews: The Search after Truth by Nicholas Malebranche; Thomas M. Lennon; Paul J. Olscamp; Elucidations of the Search after Truth by Thomas M. Lennon; Philosophical Commentary by Thomas M. Lennon. Philosophy of Science, 49(1):146–147, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186890.

Burian:1982:BRB

[157] Richard M. Burian. Book review: Problems of Genetics by William Bateson. Philosophy of Science, 49(1):147–149, March 1982. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186891.

Engelhardt:1982:BRB

[158] H. Tristram Engelhardt, Jr. Book review: The Philosophy of Medicine: The Early Eighteenth Century by Lester S. King. Philosophy of Science, 49(1):149–150, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186892.

Anonymous:1982:BMa

[159] Anonymous. Back matter. Philosophy of Science, 49(1):151–157, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186893.

Anonymous:1982:FMa

[160] Anonymous. Front matter. Philosophy of Science, 49(1):??, March 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186876.

Sober:1982:ACG

[161] Elliott Sober and Richard C. Lewontin. Artifact, cause and genic selection. Philosophy of Science, 49(2):157–180, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186916.

Niiniluoto:1982:WSW

[162] Ilkka Niiniluoto. What shall we do with verisimilitude? Philosophy of Science, 49(2):181–197, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186917.

McLaughlin:1982:IIL

[163] Robert McLaughlin. Invention and induction Laudan, Simon and the logic of discovery. Philosophy of Science, 49(2):198–211, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186918.

Robinson:1982:SM

[164] William S. Robinson. Sellarsian materialism. Philosophy of Science, 49 (2):212–227, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186919.

Aidun:1982:AFN

[165] John Aidun. Aristotelian force as Newtonian power. Philosophy of Science, 49(2):228–235, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/186920.

Amundson:1982:SEE

[166] Ron Amundson. Science, ethnoscience, and ethnocentrism. Philosophy of Science, 49(2):236–250, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186921.

Good:1982:GEE

[167] I. J. Good. A good explanation of an event is not necessarily corroborated by the event. Philosophy of Science, 49(2):251–253, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186922.

Hoy:1982:AST

[168] Ronald C. Hoy. Ambiguities in the subjective timing of experiences debate. Philosophy of Science, 49(2):254–262, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186923.

Currie:1982:NR

[169] Gregory Currie. A note on realism. Philosophy of Science, 49(2):263– 267, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186924.

Rosenberg:1982:PDF

[170] Alexander Rosenberg. On the propensity definition of fitness. Philosophy of Science, 49(2):268–273, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

186925.

Laymon:1982:ITE

[171] Ronald Laymon. Independent testability and experimental type: Response to Erlichson. Philosophy of Science, 49(2):274–281, June 1982.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186926. See [136].

Smith:1982:RFC

[172] Joseph Wayne Smith. A reply to Frankel’s criticism of Harre’s theory of causality. Philosophy of Science, 49(2):282–289, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186927.

Hanson:1982:BRB

[173] Philip P. Hanson and Edwin Levy. Book review: The Scientific Image by Bas C. van Fraassen. Philosophy of Science, 49(2):290–293, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186928.

Spielman:1982:BRB

[174] Stephen Spielman. Book review: Studies in Inductive Logic and Probability: Volume II by Richard C. Jeffrey. Philosophy of Science, 49(2): 293–295, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/186929.

Nicholas:1982:BRB

[175] John Nicholas. Book review: Black Body Theory and the Quantum Discontinuity, 1894–1912 by T. S. Kuhn. Philosophy of Science, 49(2):295– 297, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/186930.

Loewer:1982:BRB

[176] Barry Loewer. Book review: Knowledge and the Flow of Information by Fred I. Dretske. Philosophy of Science, 49(2):297–300, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/186931.

Anonymous:1982:CQL

[177] Anonymous. Corrections: Quantum logic and the projection postulate. Philosophy of Science, 49(2):301, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/186932. See [118].

Anonymous:1982:BR

[178] Anonymous. Books received. Philosophy of Science, 49(2):302–307, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/186933.

Anonymous:1982:BMb

[179] Anonymous. Back matter. Philosophy of Science, 49(2):308, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186934.

Anonymous:1982:FMb

[180] Anonymous. Front matter. Philosophy of Science, 49(2):??, June 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/186915.

Hollinger:1982:IMI

[181] Henry B. Hollinger and Michael J. Zenzen. An interpretation of macroscopic irreversibility within the Newtonian framework. Philosophy of Science, 49(3):309–354, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187279.

Bell:1982:LQL

[182] John Bell and Michael Hallett. Logic, quantum logic and empiricism. Philosophy of Science, 49(3):355–379, September 1982. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187280. See reply [477].

Pitowsky:1982:STQ

[183] Itamar Pitowsky. Substitution and truth in quantum logic. Philosophy of Science, 49(3):380–401, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187281.

Bub:1982:QLC

[184] Jeffrey Bub. Quantum logic, conditional probability, and interference. Philosophy of Science, 49(3):402–421, September 1982. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187282.

Stairs:1982:QLL

[185] Allen Stairs. Quantum logic and the Lüders Rule. Philosophy of Science, 49(3):422–436, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187283.

Mackinnon:1982:TSC

[186] Edward Mackinnon. The truth of scientific claims. Philosophy of Science, 49(3):437–462, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187284.

Spirtes:1982:STS

[187] Peter Spirtes and Clark Glymour. Space–time and synonymy. Philosophy of Science, 49(3):463–477, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187285.

Jammer:1982:NPG

[188] Max Jammer. A note on Peter Gibbins’ “A Note on Quantum Logic and the Uncertainty Principle”. Philosophy of Science, 49(3):478–479, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187286. See [85].

Bub:1982:BRB

[189] Jeffrey Bub. Book review: Readings from the New Book on Nature — Physics and Metaphysics in the Modern Novel by Robert Nadeau. Philosophy of Science, 49(3):480–481, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187287.

Stairs:1982:BRB

[190] Allen Stairs. Book review: Studies in the Foundations of Quantum Mechanics by Patrick Suppes. Philosophy of Science, 49(3):481–484, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187288.

Anonymous:1982:BMc

[191] Anonymous. Back matter. Philosophy of Science, 49(3):485–487, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187289.

Anonymous:1982:FMc

[192] Anonymous. Front matter. Philosophy of Science, 49(3):??, September 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187278.

Shapere:1982:COS

[193] Dudley Shapere. The concept of observation in science and philosophy. Philosophy of Science, 49(4):485–525, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187163.

Kourany:1982:TEA

[194] Janet A. Kourany. Towards an empirically adequate theory of science. Philosophy of Science, 49(4):526–548, December 1982. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187164.

Bechtel:1982:TCE

[195] William Bechtel. Two common errors in explaining biological and psychological phenomena. Philosophy of Science, 49(4):549–574, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187165.

Manders:1982:STO

[196] Kenneth L. Manders. On the space–time ontology of physical theories. Philosophy of Science, 49(4):575–590, December 1982. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187166.

Casullo:1982:PSI

[197] Albert Casullo. Particulars, substrata, and the identity of indiscernibles. Philosophy of Science, 49(4):591–603, December 1982. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187167.

Hardin:1982:DCR

[198] Clyde L. Hardin and Alexander Rosenberg. In defense of convergent realism. Philosophy of Science, 49(4):604–615, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187168.

Woller:1982:HMM

[199] Raymond Woller. Harre and Madden’s multifarious account of natural necessity. Philosophy of Science, 49(4):616–632, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187169.

Kitcher:1982:GRF

[200] Patricia Kitcher. Genetics, reduction and functional psychology. Philosophy of Science, 49(4):633–636, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187170.

Rozeboom:1982:LDH

[201] William W. Rozeboom. Let’s dump hypothetico-deductivism for the right reasons. Philosophy of Science, 49(4):637–647, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187171.

Stabler:1982:BRB

[202] Edward P. Stabler, Jr. Book review: Sociobiology and the Preemption of Social Science by Alexander Rosenberg. Philosophy of Science, 49(4): 648–651, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187172.

McMillan:1982:BRB

[203] John McMillan. Book review: Capital, Profits and Prices: An Essay in the Philosophy of Economics by Daniel M. Hausman. Philosophy of Science, 49(4):651–653, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187173.

Nickles:1982:BRB

[204] Thomas Nickles. Book review: Science and Hypothesis by Larry Laudan. Philosophy of Science, 49(4):653–655, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187174.

Brush:1982:BRB

[205] Stephen G. Brush. Book review: Conceptions of Ether. Studies in the History of Ether Theories 1740–1900 by G. N. Cantor; M. J. S. Hodge. Philosophy of Science, 49(4):655–656, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187175.

Anonymous:1982:BMd

[206] Anonymous. Back matter. Philosophy of Science, 49(4):657–666, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187176.

Anonymous:1982:VI

[207] Anonymous. Volume information. Philosophy of Science, 49(4):659–663, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187161.

Anonymous:1982:FMd

[208] Anonymous. Front matter. Philosophy of Science, 49(4):??, December 1982. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187162.

Hanna:1983:EA

[209] Joseph F. Hanna. Empirical adequacy. Philosophy of Science, 50(1):1– 34, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187402.

Eells:1983:PCQ

[210] Ellery Eells and Elliott Sober. Probabilistic causality and the question of transitivity. Philosophy of Science, 50(1):35–57, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187403.

Hausman:1983:TCR

[211] Daniel Hausman. Are there causal relations among dependent variables? Philosophy of Science, 50(1):58–81, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187404.

Weirich:1983:CPP

[212] Paul Weirich. Conditional probabilities and probabilities given knowledge of a condition. Philosophy of Science, 50(1):82–95, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187405.

Cummins:1983:ASB

[213] Robert Cummins. Analysis and subsumption in the behaviorism of Hull. Philosophy of Science, 50(1):96–111, March 1983. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187406.

Lloyd:1983:NDS

[214] Elisabeth A. Lloyd. The nature of Darwin’s support for the theory of natural selection. Philosophy of Science, 50(1):112–129, March 1983. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187407.

Acock:1983:AU

[215] Malcolm Acock. The age of the universe. Philosophy of Science, 50(1): 130–145, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187408.

Schmaus:1983:DHL

[216] Warren Schmaus. In defense of historical laws. Philosophy of Science, 50(1):146–150, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187409.

Nissen:1983:WTD

[217] Lowell Nissen. Wright on teleological descriptions of goal-directed behavior. Philosophy of Science, 50(1):151–158, March 1983. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187410.

Kroes:1983:CPH

[218] Peter Kroes. The clock paradox, or how to get rid of absolute time. Philosophy of Science, 50(1):159–163, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187411.

Katz:1983:CI

[219] Jonathan Katz. Causality and indeterminism. Philosophy of Science, 50 (1):164–166, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187412.

Horwich:1983:BRB

[220] Paul Horwich. Book review: Applications of Inductive Logic by L. Jonathan Cohen; Mary Hesse. Philosophy of Science, 50(1):167–168, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187413.

Enfield:1983:BRB

[221] Patrick Enfield. Book review: Reduction, Time and Reality by Richard Healey. Philosophy of Science, 50(1):168–169, March 1983. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187414.

Airaksinen:1983:BRB

[222] Timo Airaksinen. Book review: The Philosophy of Nicholas Rescher: Discussion and Replies by Ernest Sosa. Philosophy of Science, 50(1):169– 171, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187415.

McMullin:1983:BRB

[223] Ernan McMullin. Book review: Prelude to Galileo by William A. Wallace. Philosophy of Science, 50(1):171–173, March 1983. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187416.

Marras:1983:BRB

[224] Ausonio Marras. Book review: Language and Learning: The Debate between Jean Piaget and Noam Chomsky by Massimo Piattelli-Palmarini. Philosophy of Science, 50(1):173–175, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/187417.

Quinton:1983:BRB

[225] Reg Quinton. Book review: Readings in the Philosophy of Psychology, Vol. 2 by Ned Block. Philosophy of Science, 50(1):175–176, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187418.

Anonymous:1983:BMa

[226] Anonymous. Back matter. Philosophy of Science, 50(1):177–179, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187419.

Anonymous:1983:FMa

[227] Anonymous. Front matter. Philosophy of Science, 50(1):??, March 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187401.

Heelan:1983:NSH

[228] Patrick Heelan. Natural science as a hermeneutic of instrumentation. Philosophy of Science, 50(2):181–204, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188010.

Mundy:1983:RTE

[229] Brent Mundy. Relational theories of Euclidean space and Minkowski spacetime. Philosophy of Science, 50(2):205–226, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188011.

Hellman:1983:RP

[230] Geoffrey Hellman. Realist principles. Philosophy of Science, 50(2):227– 249, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188012.

Thagard:1983:RC

[231] Paul Thagard and Richard E. Nisbett. Rationality and charity. Philosophy of Science, 50(2):250–267, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188013.

Amundson:1983:CTI

[232] Ron Amundson. E. C. Tolman and the intervening variable: A study in the epistemological history of psychology. Philosophy of Science, 50(2): 268–282, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188014.

Good:1983:PED

[233] I. J. Good. The philosophy of exploratory data analysis. Philosophy of Science, 50(2):283–295, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188015.

Gibson:1983:T

[234] Quentin Gibson. Tendencies. Philosophy of Science, 50(2):296–308, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188016.

Teller:1983:QPI

[235] Paul Teller. Quantum physics, the identity of indiscernibles, and some unanswered questions. Philosophy of Science, 50(2):309–319, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188017.

Laudan:1983:IJ

[236] Larry Laudan. Invention and justification. Philosophy of Science, 50(2): 320–322, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188018.

Brown:1983:BR

[237] Harold I. Brown. Being rational. Philosophy of Science, 50(2):323–325, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188019.

Conee:1983:SNJ

[238] Earl Conee and Richard Feldman. Stich and Nisbett on justifying inference rules. Philosophy of Science, 50(2):326–331, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188020.

Rosenberg:1983:CEG

[239] Alexander Rosenberg. Coefficients, effects, and genic selection. Philosophy of Science, 50(2):332–338, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

188021.

Matthews:1983:BRB

[240] Robert J. Matthews. Book review: Inquiries and Provocations: Selected Writings, 1929–1974 by Herbert Feigl. Philosophy of Science, 50(2):339– 344, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188022.

Martin:1983:BRB

[241] R. Naill D. Martin. Book review: Science and Society: Studies in the Sociology of Science by Joseph Agassi; Robert S. Cohen; Marx W. Wartofsky. Philosophy of Science, 50(2):345–346, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188023.

Duncan:1983:BRBa

[242] Howard Duncan. Book review: Space and Incongruence: The Origin of Kant’s Idealism by Jill Vance Buroker. Philosophy of Science, 50(2):346– 348, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/188024.

Olson:1983:BRB

[243] Richard G. Olson. Book review: Sir William Rowan Hamilton by Thomas L. Hankins. Philosophy of Science, 50(2):348–349, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188025.

Beck:1983:BRB

[244] Lewis White Beck. Book review: Enzyklopadie Philosophie und Wissenschaftstheorie by Jurgen Mittelstrass. Philosophy of Science, 50(2): 349–350, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188026.

Lugg:1983:BRB

[245] Andrew Lugg. Book review: Popper and beyond by David Stove. Philosophy of Science, 50(2):350–352, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188027.

Anonymous:1983:BR

[246] Anonymous. Books received. Philosophy of Science, 50(2):353–356, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188028.

Anonymous:1983:BMb

[247] Anonymous. Back matter. Philosophy of Science, 50(2):357, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188029.

Anonymous:1983:FMb

[248] Anonymous. Front matter. Philosophy of Science, 50(2):??, June 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188009.

Mermin:1983:PDC

[249] N. D. Mermin. Pair distributions and conditional independence: Some hints about the structure of strange quantum correlations. Philosophy of Science, 50(3):359–373, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187854.

Kyburg:1983:RC

[250] Henry E. Kyburg, Jr. The reference class. Philosophy of Science, 50(3): 374–397, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187855.

Seidenfeld:1983:CBF

[251] Teddy Seidenfeld and Mark J. Schervish. A conflict between finite additivity and avoiding Dutch book. Philosophy of Science, 50(3):398–412, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187856.

Teller:1983:PPF

[252] Paul Teller. The projection postulate as a fortuitous approximation. Philosophy of Science, 50(3):413–431, September 1983. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187857.

Thompson:1983:TME

[253] Paul Thompson. Tempo and mode in evolution: Punctuated equilibria and the modern synthetic theory. Philosophy of Science, 50(3):432–452, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187858.

Madan:1983:ITS

[254] Dilip B. Madan. Inconsistent theories as scientific objectives. Philosophy of Science, 50(3):453–470, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187859.

Christensen:1983:GER

[255] David Christensen. Glymour on evidential relevance. Philosophy of Science, 50(3):471–481, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187860.

Roth:1983:SNE

[256] Paul A. Roth. Siegel on naturalized epistemology and natural science. Philosophy of Science, 50(3):482–493, September 1983. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187861.

Galle:1983:KPO

[257] Peter Galle. Kordig’s paradox objection to radical meaning variance theories. Philosophy of Science, 50(3):494–497, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187862.

Friedman:1983:BRB

[258] Michael Friedman. Book review: Philosophical Papers by Moritz Schlick; H. L. Mulder; B. F. B. van de Velde-Schlick. Philosophy of Science, 50(3): 498–514, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187863.

Bub:1983:BRB

[259] Jeffrey Bub. Book review: Physics and Philosophy: Selected Essays by Henry Margenau. Philosophy of Science, 50(3):515–516, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187864.

Arthur:1983:BRB

[260] Ric Arthur. Book review: The Investigation of the Physical World by G. Toraldo Di Francia. Philosophy of Science, 50(3):516–518, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187865.

Mundy:1983:BRB

[261] Brent Mundy. Book review: Understanding the Space–time Concepts of Special Relativity by Arthur Evett. Philosophy of Science, 50(3):518, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187866.

Nute:1983:BRB

[262] Donald Nute. Book review: Ifs: Conditionals, Belief, Decision, Chance, and Time by William L. Harper; Robert Stalnaker; Glenn Pearce. Philosophy of Science, 50(3):518–520, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187867.

Anonymous:1983:BMc

[263] Anonymous. Back matter. Philosophy of Science, 50(3):521–522, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187868.

Anonymous:1983:FMc

[264] Anonymous. Front matter. Philosophy of Science, 50(3):??, September 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187853.

Shapiro:1983:MR

[265] Stewart Shapiro. Mathematics and reality. Philosophy of Science, 50(4): 523–548, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187555.

Kaplan:1983:DTP

[266] Mark Kaplan. Decision theory as philosophy. Philosophy of Science, 50 (4):549–577, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187556.

Stairs:1983:QLR

[267] Allen Stairs. Quantum logic, realism, and value definiteness. Philosophy of Science, 50(4):578–602, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187557.

Rohrlich:1983:ET

[268] Fritz Rohrlich and Larry Hardin. Established theories. Philosophy of Science, 50(4):603–617, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187558.

Collier:1983:FDC

[269] John Collier. Frequency-dependent causation: A defense of Giere. Philosophy of Science, 50(4):618–625, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187559.

Glymour:1983:RBT

[270] Clark Glymour. Revisions of bootstrap testing. Philosophy of Science, 50 (4):626–629, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187560.

Kyburg:1983:LPD

[271] Henry E. Kyburg, Jr. Levi, Petersen, and direct inference. Philosophy of Science, 50(4):630–634, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187561.

Levi:1983:KRD

[272] Isaac Levi. Kyburg on random designators. Philosophy of Science, 50 (4):635–642, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187562.

Robinson:1983:IET

[273] Joseph D. Robinson. Identifying error types on behalf of better science. Philosophy of Science, 50(4):643–647, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187563.

Sober:1983:RRG

[274] Elliott Sober and Richard C. Lewontin. Reply to Rosenberg on genic selectionism. Philosophy of Science, 50(4):648–650, December 1983. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187564.

Mermin:1983:CNB

[275] N. D. Mermin. Critical notice: The Open Universe by Karl R. Popper; Quantum Theory and the Schism in Physics by Karl R. Popper; the great quantum muddle. Philosophy of Science, 50(4):651–656, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187565.

Bunn:1983:BRB

[276] Robert Bunn. Book review: Revolutions & Reconstructions in the Philosophy of Science by Mary Hesse. Philosophy of Science, 50(4):657–659, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187566.

Morton:1983:BRB

[277] Adam Morton. Book review: Probability and Evidence by Paul Horwich. Philosophy of Science, 50(4):659–660, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187567.

Shrader:1983:BRB

[278] Douglas Shrader. Book review: Scientific Knowledge: Causation, Explanation, and Corroboration by James H. Fetzer. Philosophy of Science, 50 (4):660–662, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187568.

Shea:1983:BRB

[279] William R. Shea. Book review: Apologia del Metodo by Marcello Pera. Philosophy of Science, 50(4):662–663, December 1983. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187569.

Chandler:1983:BRB

[280] Marthe Chandler. Book review: Real Time by D. H. Mellor. Philosophy of Science, 50(4):663–665, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187570.

Boylan:1983:BRB

[281] Michael Boylan. Book reviews: Philosophy of Mathematics and Deductive Structure in Euclid’s Elements by Ian Mueller; The Beginnings of Greek Mathematics by Arpad Szabo; A. M. Ungar. Philosophy of Science, 50 (4):665–668, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187571.

Duncan:1983:BRBb

[282] Howard Duncan. Book review: Metaphysics and Natural Philosophy by P. M. Harman. Philosophy of Science, 50(4):668–669, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187572.

Healey:1983:BRB

[283] Richard A. Healey. Book review: Realism and the Aim of Science by Karl R. Popper; W. W. Bartley. Philosophy of Science, 50(4):669–671, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187573.

Robinson:1983:BRB

[284] William S. Robinson. Book review: Pictures, Images and Conceptual Change, An Analysis of Wilfrid Sellars’ Philosophy of Science by Joseph C. Pitt. Philosophy of Science, 50(4):671–672, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187574.

Winston:1983:BRB

[285] Morton Winston. Book review: The Philosophy of W. V. Quine: An Expository Essay by Roger F. Gibson. Philosophy of Science, 50(4):673– 674, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187575.

Anonymous:1983:BMd

[286] Anonymous. Back matter. Philosophy of Science, 50(4):675–686, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187576.

Anonymous:1983:VI

[287] Anonymous. Volume information. Philosophy of Science, 50(4):677–682, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187553.

Anonymous:1983:FMd

[288] Anonymous. Front matter. Philosophy of Science, 50(4):??, December 1983. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187554.

Butts:1984:PS

[289] Robert E. Butts. Philosophy of science: 1934–1984. Philosophy of Science, 51(1):1–3, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187726.

Carnap:1984:CPP

[290] Rudolf Carnap. On the character of philosophic problems. Philosophy of Science, 51(1):5–19, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187727.

Churchman:1984:EYP

[291] C. West Churchman. Early years of the philosophy of science association. Philosophy of Science, 51(1):20–22, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187728.

Fodor:1984:OR

[292] Jerry Fodor. Observation reconsidered. Philosophy of Science, 51(1):23– 43, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187729.

Klee:1984:MDC

[293] Robert L. Klee. Micro-determinism and concepts of emergence. Philosophy of Science, 51(1):44–63, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187730.

Stabler:1984:RNE

[294] Edward P. Stabler, Jr. Rationality in naturalized epistemology. Philosophy of Science, 51(1):64–78, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187731.

Foss:1984:AVF

[295] Jeff Foss. On accepting van Fraassen’s image of science. Philosophy of Science, 51(1):79–92, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187732.

Duncan:1984:ICM

[296] Howard Duncan. Inertia, the communication of motion, and Kant’s third law of mechanics. Philosophy of Science, 51(1):93–119, March 1984. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187733.

Good:1984:PME

[297] I. J. Good and Alan F. McMichael. A pragmatic modification of explicativity for the acceptance of hypotheses. Philosophy of Science, 51 (1):120–127, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187734.

Nathan:1984:FE

[298] Amos Nathan. False expectations. Philosophy of Science, 51(1):128–136, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187735.

Weitzenfeld:1984:VRA

[299] Julian S. Weitzenfeld. Valid reasoning by analogy. Philosophy of Science, 51(1):137–149, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187736.

Weingard:1984:GUG

[300] Robert Weingard. Grand unified gauge theories and the number of elementary particles. Philosophy of Science, 51(1):150–155, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187737.

Laudan:1984:RR

[301] Larry Laudan. Realism without the real. Philosophy of Science, 51(1): 156–162, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187738.

Horwich:1984:BRB

[302] Paul Horwich. Book review: Wittgenstein on Rules and Private Language by Saul Kripke. Philosophy of Science, 51(1):163–171, March 1984. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187739.

Stein:1984:BRB

[303] Howard Stein. Book review: Electricity in the 17th and 18th Centuries: A Study of Early Modern Physics by J. L. Heilbron. Philosophy of Science, 51(1):172–175, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187740.

Dillinger:1984:BRB

[304] Mike Dillinger. Book review: Language and Other Abstract Objects by J. J. Katz. Philosophy of Science, 51(1):175–176, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187741.

Jones:1984:BRBa

[305] Gary E. Jones. Book review: Induction by Nicholas Rescher. Philosophy of Science, 51(1):176–177, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187742.

Lavin:1984:BRB

[306] Michael Lavin. Book review: The Standing of Psychoanalysis by B. A. Farrell. Philosophy of Science, 51(1):177–179, March 1984. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187743.

Anonymous:1984:BMa

[307] Anonymous. Back matter. Philosophy of Science, 51(1):180–182, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187744.

Anonymous:1984:FMa

[308] Anonymous. Front matter. Philosophy of Science, 51(1):??, March 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187725.

Beatty:1984:CNS

[309] John Beatty. Chance and natural selection. Philosophy of Science, 51(2): 183–211, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187421.

Sober:1984:CCE

[310] Elliott Sober. Common cause explanation. Philosophy of Science, 51(2): 212–241, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187422.

Lloyd:1984:SAS

[311] Elisabeth A. Lloyd. A semantic approach to the structure of population genetics. Philosophy of Science, 51(2):242–264, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187423.

Bechtel:1984:RIC

[312] William Bechtel. Reconceptualizations and interfield connections: The discovery of the link between vitamins and coenzymes. Philosophy of Science, 51(2):265–292, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187424.

Holsinger:1984:NBS

[313] Kent E. Holsinger. The nature of biological species. Philosophy of Science, 51(2):293–307, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187425.

Kitcher:1984:S

[314] Philip Kitcher. Species. Philosophy of Science, 51(2):308–333, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187426. See [315].

Sober:1984:SSE

[315] Elliott Sober. Sets, species, and evolution: Comments on Philip Kitcher’s “species”. Philosophy of Science, 51(2):334–341, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187427. See [314].

Brandon:1984:PIF

[316] Robert Brandon and John Beatty. The propensity interpretation of ‘fitness’ — no interpretation is no substitute. Philosophy of Science, 51(2): 342–347, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187428.

Ruse:1984:BRB

[317] Michael Ruse. Book review: Abusing Science: The Case against Creationism by Philip Kitcher. Philosophy of Science, 51(2):348–354, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187429.

Hooker:1984:BRB

[318] Cliff Hooker. Book review: From Being to Becoming: Time and Complexity in the Physical Sciences by Ilya Prigogine. Philosophy of Science, 51(2):355–357, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187430.

Richardson:1984:BRB

[319] Robert C. Richardson. Book review: The Extended Phenotype: The Gene as the Unit of Selection by Richard Dawkins. Philosophy of Science, 51 (2):357–359, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187431.

Sarkar:1984:BRB

[320] Sahotra Sarkar. Book review: Hierarchy: Perspectives for Ecological Complexity by T. F. H. Allen; Thomas B. Starr. Philosophy of Science, 51(2): 359–361, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187432.

Anonymous:1984:BR

[321] Anonymous. Books received. Philosophy of Science, 51(2):362–365, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187433.

Anonymous:1984:BMb

[322] Anonymous. Back matter. Philosophy of Science, 51(2):366–367, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187434.

Anonymous:1984:FMb

[323] Anonymous. Front matter. Philosophy of Science, 51(2):??, June 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187420.

Teller:1984:PPN

[324] Paul Teller. The projection postulate: A new perspective. Philosophy of Science, 51(3):369–395, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187490.

Richardson:1984:BII

[325] Robert C. Richardson. Biology and ideology: The interpenetration of science and values. Philosophy of Science, 51(3):396–420, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187491.

Woodward:1984:EA

[326] James Woodward. Explanatory asymmetries. Philosophy of Science, 51 (3):421–442, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187492.

Forge:1984:TFT

[327] John Forge. Theoretical functions, theory and evidence. Philosophy of Science, 51(3):443–463, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187493.

Peterson:1984:SIS

[328] Philip L. Peterson. Semantic indeterminacy and scientific underdetermination. Philosophy of Science, 51(3):464–487, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187494.

Nissen:1984:WAT

[329] Lowell Nissen. Woodfield’s analysis of teleology. Philosophy of Science, 51 (3):488–494, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187495.

Hands:1984:WEE

[330] Douglas W. Hands. What economics is not: An Economist’s response to Rosenberg. Philosophy of Science, 51(3):495–503, September 1984.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187496.

Wallace:1984:GCT

[331] William A. Wallace. Galileo and the continuity thesis. Philosophy of Science, 51(3):504–510, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187497.

Smith:1984:WOF

[332] Joseph Wayne Smith and Sharyn Ward. Are we only five minutes old? Acock on the age of the universe. Philosophy of Science, 51(3):511–513, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187498.

Hochberg:1984:BRB

[333] Herbert Hochberg. Book review: Concept and Quality by George Bealer. Philosophy of Science, 51(3):514–516, September 1984. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187499.

Fetzer:1984:BRB

[334] James H. Fetzer. Book review: The Nature of Explanation by Peter Achinstein. Philosophy of Science, 51(3):516–519, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187500.

Jones:1984:BRBb

[335] Roger Jones. Book review: From Calculus to Set Theory, 1630–1910: An Introductory History by I. Grattan-Guinness. Philosophy of Science, 51 (3):519–522, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187501.

Bantz:1984:BRB

[336] David A. Bantz. Book review: Limits of Scientific Inquiry by Gerald Holton; Robert S. Morison. Philosophy of Science, 51(3):522–525, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187502.

Sylla:1984:BRB

[337] Edith Dudley Sylla. Book review: Theory Change, Ancient Axiomatics, and Galileo’s Methodology by J. Hintikka; D. Gruender; E. Agazzi. Philosophy of Science, 51(3):525–527, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187503.

Anonymous:1984:BMc

[338] Anonymous. Back matter. Philosophy of Science, 51(3):??, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187504.

Anonymous:1984:FMc

[339] Anonymous. Front matter. Philosophy of Science, 51(3):??, September 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187489.

Horgan:1984:SN

[340] Terence Horgan. Science nominalized. Philosophy of Science, 51(4):529– 549, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187974.

Papineau:1984:RE

[341] David Papineau. Representation and explanation. Philosophy of Science, 51(4):550–572, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187975.

Nelson:1984:SIS

[342] Alan Nelson. Some issues surrounding the reduction of macroeconomics to microeconomics. Philosophy of Science, 51(4):573–594, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187976.

Rothbart:1984:SMS

[343] Daniel Rothbart. The semantics of metaphor and the structure of science. Philosophy of Science, 51(4):595–615, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187977.

Kitcher:1984:AMM

[344] Philip Kitcher. Against the monism of the moment: A reply to Elliott Sober. Philosophy of Science, 51(4):616–630, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187978.

Bonevac:1984:SSS

[345] Daniel Bonevac. Systems of substitutional semantics. Philosophy of Science, 51(4):631–656, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187979.

Amundson:1984:CHR

[346] Ron Amundson and Laurence D. Smith. Clark hull, Robert cummins, and functional analysis. Philosophy of Science, 51(4):657–666, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187980.

Siegel:1984:EPN

[347] Harvey Siegel. Empirical psychology, naturalized epistemology, and first philosophy. Philosophy of Science, 51(4):667–676, December 1984. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187981.

Nathan:1984:FID

[348] Amos Nathan. The fallacy of intrinsic distributions. Philosophy of Science, 51(4):677–684, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187982.

Pitowsky:1984:UFT

[349] Itamar Pitowsky. Unified field theory and the conventionality of geometry. Philosophy of Science, 51(4):685–689, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187983.

Maher:1984:BRB

[350] Patrick Maher. Book review: The Enterprise of Knowledge: An Essay on Knowledge, Credal Probability, and Chance by Isaac Levi. Philosophy of Science, 51(4):690–692, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187984.

Okruhlik:1984:BRB

[351] Kathleen Okruhlik. Book review: Reason, Truth and History by Hilary Putnam. Philosophy of Science, 51(4):692–694, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187985.

Hoy:1984:BRB

[352] Ronald C. Hoy. Book review: Time: A Philosophical Analysis by T. Chapman. Philosophy of Science, 51(4):694–696, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187986.

Smith:1984:BRB

[353] Terry L. Smith. Book review: Behaviorism, Science, and Human Nature by Barry Schwartz; Hugh Lacey. Philosophy of Science, 51(4):696–698, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187987.

Riepe:1984:BRB

[354] Dale Riepe. Book review: Studies in the History of Science in India by Debiprasad Chattopadhyaya. Philosophy of Science, 51(4):698–699, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187988.

Anonymous:1984:BMd

[355] Anonymous. Back matter. Philosophy of Science, 51(4):701–712, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187989.

Anonymous:1984:VI

[356] Anonymous. Volume information. Philosophy of Science, 51(4):703–707, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187972.

Anonymous:1984:FMd

[357] Anonymous. Front matter. Philosophy of Science, 51(4):??, December 1984. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187973.

Kyburg:1985:CQL

[358] Henry E. Kyburg, Jr. The confirmation of quantitative laws. Philosophy of Science, 52(1):1–22, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187595.

Maxwell:1985:PSR

[359] Nicholas Maxwell. Are probabilism and Special Relativity incompatible? Philosophy of Science, 52(1):23–43, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187596.

Nugayev:1985:HQM

[360] R. M. Nugayev. The history of quantum mechanics as a decisive argument favoring Einstein over Lorentz. Philosophy of Science, 52(1):44–63, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187597.

Busch:1985:NQT

[361] Paul Busch and Pekka J. Lahti. A note on quantum theory, complementarity, and uncertainty. Philosophy of Science, 52(1):64–77, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187598.

Kitcher:1985:NTW

[362] Patricia Kitcher. Narrow taxonomy and Wide functionalism. Philosophy of Science, 52(1):78–97, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187599.

Stidd:1985:PWA

[363] Benton M. Stidd. Are punctuationists wrong about the modern synthesis? Philosophy of Science, 52(1):98–109, March 1985. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187600.

Otte:1985:PCS

[364] Richard Otte. Probabilistic causality and Simpson’s paradox. Philosophy of Science, 52(1):110–125, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187601.

Pearce:1985:AED

[365] David Pearce and Veikko Rantala. Approximative explanation is deductive-nomological. Philosophy of Science, 52(1):126–140, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187602.

Sober:1985:WWH

[366] Elliott Sober. What would happen if everyone did it? A reply to Collier and Giere on frequency dependent causation. Philosophy of Science, 52 (1):141–150, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187603.

See response [534].

Thomason:1985:NTS

[367] Richmond H. Thomason. Note on tense and subjunctive conditionals. Philosophy of Science, 52(1):151–153, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187604.

Pitowsky:1985:QMV

[368] Itamar Pitowsky. Quantum mechanics and value definiteness. Philosophy of Science, 52(1):154–156, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187605.

Stabler:1985:BRB

[369] Edward P. Stabler, Jr. Book review: Mind and Meaning by Brian Loar. Philosophy of Science, 52(1):157–159, March 1985. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187606.

Brown:1985:BRBa

[370] Harold I. Brown. Book review: Space-Perception and the Philosophy of Science by Patrick A. Heelan. Philosophy of Science, 52(1):159–160, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187607.

Holsinger:1985:BRB

[371] Kent E. Holsinger. Book review: Dimensions of Darwinism: Themes & Counterthemes in Twentieth-Century Evolutionary Theory by Marjorie Grene. Philosophy of Science, 52(1):161–163, March 1985. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187608.

Burian:1985:BRB

[372] Richard M. Burian. Book review: Radical Knowledge: A Philosophical Inquiry into the Nature and Limits of Science by Gonzalo Munévar. Philosophy of Science, 52(1):163–165, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187609.

Usselman:1985:BRB

[373] M. C. Usselman. Book review: The Formation of the German Chemical Community (1720–1795) by Karl Hufbauer. Philosophy of Science, 52 (1):165–166, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187610.

Mundy:1985:BRBa

[374] Brent Mundy. Book review: Polish Essays in the Philosophy of the Natural Sciences by Wladyslaw Krajewski. Philosophy of Science, 52(1):166–167, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187611.

Brown:1985:BRBb

[375] James Robert Brown. Book review: Science and Convention by Jerzy Giedymin. Philosophy of Science, 52(1):168–169, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187612.

Gudder:1985:BRB

[376] Stanley P. Gudder. Book review: The Wave-Particle Dualism: A Tribute to Louis de Broglie on His 90th Birthday by S. Diner; D. Fargue; G. Lochak; F. Selleri. Philosophy of Science, 52(1):169–170, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187613.

Michalos:1985:BRB

[377] Alex C. Michalos. Book review: Epistemology, Methodology, and the Social Sciences by Robert S. Cohen; Marx W. Wartofsky. Philosophy of Science, 52(1):170–171, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187614.

Anonymous:1985:BMa

[378] Anonymous. Back matter. Philosophy of Science, 52(1):172–176, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187615.

Anonymous:1985:FMa

[379] Anonymous. Front matter. Philosophy of Science, 52(1):??, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187594.

Nickles:1985:BDC

[380] Thomas Nickles. Beyond divorce: Current status of the discovery debate. Philosophy of Science, 52(2):177–206, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187506.

Lugg:1985:PD

[381] Andrew Lugg. The process of discovery. Philosophy of Science, 52(2):207– 220, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187507.

Resnik:1985:LND

[382] Michael D. Resnik. Logic: Normative or descriptive? The ethics of belief or a branch of psychology? Philosophy of Science, 52(2):221–238, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187508.

Fisch:1985:WCI

[383] Menachem Fisch. Whewell’s consilience of inductions — an evaluation. Philosophy of Science, 52(2):239–255, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187509.

Urbach:1985:RDE

[384] Peter Urbach. Randomization and the design of experiments. Philosophy of Science, 52(2):256–273, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187510. See comments [94, 571].

Seidenfeld:1985:CCS

[385] Teddy Seidenfeld. Calibration, coherence, and scoring rules. Philosophy of Science, 52(2):274–294, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187511.

Marras:1985:CMS

[386] Ausonio Marras. The Churchlands on methodological solipsism and computational psychology. Philosophy of Science, 52(2):295–309, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187512.

Nickles:1985:BRB

[387] Thomas Nickles. Book review: Reason and the Search for Knowledge by Dudley Shapere. Philosophy of Science, 52(2):310–312, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187513.

Maddy:1985:BRB

[388] Penelope Maddy. Book review: The Nature of Mathematical Knowledge by Philip Kitcher. Philosophy of Science, 52(2):312–314, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187514.

Leplin:1985:BRBa

[389] Jarrett Leplin. Book review: Representing and Intervening: Introductory Topics in the Philosophy of Natural Science by Ian Hacking. Philosophy of Science, 52(2):314–315, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187515.

Garber:1985:BRB

[390] Daniel Garber. Book review: A Theory of Method by Husain Sarkar. Philosophy of Science, 52(2):315–317, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187516.

Brown:1985:BRBc

[391] James Robert Brown. Book review: Otto Hahn and the Rise of Nuclear Physics by W. Shea. Philosophy of Science, 52(2):317–318, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187517.

Mundy:1985:BRBb

[392] Brent Mundy. Book review: Physics, Philosophy, and Psychoanalysis by R. S. Cohen; L. Laudan. Philosophy of Science, 52(2):318–320, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187518.

Lavin:1985:BRB

[393] Michael Lavin. Book review: Mind and Medicine: Problems of Explanation and Evaluation in Psychiatry and the Biomedical Sciences by Larry Laudan. Philosophy of Science, 52(2):321–323, June 1985. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187519.

Grontkowski:1985:BRB

[394] C. R. Grontkowski. Book review: A Feeling for the Organism: The Life and Work of Barbara McClintock by Evelyn Fox Keller. Philosophy of Science, 52(2):323–324, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187520.

Anonymous:1985:BR

[395] Anonymous. Books received. Philosophy of Science, 52(2):325–328, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187521.

Anonymous:1985:BMb

[396] Anonymous. Back matter. Philosophy of Science, 52(2):329–330, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187522.

Anonymous:1985:FMb

[397] Anonymous. Front matter. Philosophy of Science, 52(2):??, June 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187505.

Giere:1985:PSN

[398] Ronald N. Giere. Philosophy of science naturalized. Philosophy of Science, 52(3):331–356, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187707.

Jobe:1985:ECC

[399] Evan K. Jobe. Explanation, causality, and counterfactuals. Philosophy of Science, 52(3):357–389, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187708.

Levi:1985:IIP

[400] Isaac Levi. Imprecision and indeterminacy in probability judgment. Philosophy of Science, 52(3):390–409, September 1985. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187709.

Mulaik:1985:ESE

[401] Stanley A. Mulaik. Exploratory statistics and empiricism. Philosophy of Science, 52(3):410–430, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187710.

Horwich:1985:DTL

[402] Paul Horwich. Decision theory in light of Newcomb’s problem. Philosophy of Science, 52(3):431–450, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187711.

Maloney:1985:MSR

[403] J. Christopher Maloney. Methodological solipsism reconsidered as a research strategy in cognitive psychology. Philosophy of Science, 52(3): 451–469, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187712.

Sudbery:1985:PVE

[404] A. Sudbery. Popper’s variant of the EPR experiment does not test the Copenhagen interpretation. Philosophy of Science, 52(3):470–476, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187713.

Bach:1985:BRB

[405] Kent Bach. Book review: Language, Thought, and Other Biological Categories: New Foundations for Realism by Ruth Garrett Millikan. Philosophy of Science, 52(3):477–478, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187714.

Forster:1985:BRB

[406] Malcolm R. Forster. Book review: How the Laws of Physics Lie by Nancy Cartwright. Philosophy of Science, 52(3):478–480, September 1985. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187715.

Cartwright:1985:BRB

[407] Nancy Cartwright. Book review: Quantum Theory and Measurement by John Archibald Wheeler; Wojciech Hubert Zurek. Philosophy of Science, 52(3):480–481, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187716.

Franklin:1985:BRB

[408] Allan Franklin. Book review: Scientific Explanation and Atomic Physics by Edward M. MacKinnon. Philosophy of Science, 52(3):481–483, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187717.

Harre:1985:BRB

[409] Rom Harre. Book review: A Realist Philosophy of Science by Jerrold L. Aronson. Philosophy of Science, 52(3):483–485, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187718.

Leslie:1985:BRB

[410] John Leslie. Book review: The Riddle of Existence: An Essay in Idealistic Metaphysics by Nicholas Rescher. Philosophy of Science, 52(3):485–486, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187719.

Shea:1985:BRBa

[411] William R. Shea. Book review: Hegel and the Sciences by Robert S. Cohen; Marx W. Wartofsky. Philosophy of Science, 52(3):487, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187720.

Shea:1985:BRBb

[412] William R. Shea. Book review: Grundriss der Geschichte der Philosophie, Die Philosophie der Antike, Band 3: Aeltere Akademie, Aristoteles, Peripatos. Philosophy of Science, 52(3):487–488, September 1985. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187721.

Shapiro:1985:BRB

[413] Stewart Shapiro. Book review: Philosophy of Mathematics by Paul Benacerraf; Hilary Putnam. Philosophy of Science, 52(3):488–489, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187722.

Jarvie:1985:BRB

[414] I. C. Jarvie. Book review: Principles of Scientific Sociology by Walter L. Wallace. Philosophy of Science, 52(3):489–491, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187723.

Anonymous:1985:BMc

[415] Anonymous. Back matter. Philosophy of Science, 52(3):492, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187724.

Anonymous:1985:FMc

[416] Anonymous. Front matter. Philosophy of Science, 52(3):??, September 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187706.

Mayo:1985:BEL

[417] Deborah G. Mayo. Behavioristic, evidentialist, and learning models of statistical testing. Philosophy of Science, 52(4):493–516, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187437.

Siegel:1985:WQC

[418] Harvey Siegel. What is the question concerning the rationality of science? Philosophy of Science, 52(4):517–537, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187438.

Sharp:1985:FPM

[419] W. D. Sharp and N. Shanks. Fine’s prism models for quantum correlation statistics. Philosophy of Science, 52(4):538–564, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187439.

Ringen:1985:OCP

[420] Jon Ringen. Operant conditioning and a paradox of teleology. Philosophy of Science, 52(4):565–577, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187440.

Williams:1985:SIT

[421] Mary B. Williams. Species are individuals: Theoretical foundations for the claim. Philosophy of Science, 52(4):578–590, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187441. See reply [617].

Wilson:1985:DDH

[422] Fred Wilson. Dispositions defined: Harré and Madden on analyzing disposition concepts. Philosophy of Science, 52(4):591–607, December 1985.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187442.

Skyrms:1985:UPC

[423] Brian Skyrms. Ultimate and proximate consequences in causal decision theory. Philosophy of Science, 52(4):608–611, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187443.

Miller:1985:CGT

[424] Gene Miller. Correlations and Giere’s theory of causation. Philosophy of Science, 52(4):612–614, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187444.

Malament:1985:MRA

[425] David Malament. A Modest remark about Reichenbach, rotation, and General Relativity. Philosophy of Science, 52(4):615–620, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187445.

Demopoulos:1985:BRA

[426] William Demopoulos and Michael Friedman. Bertrand Russell’s the analysis of matter: Its historical context and contemporary interest. Philosophy of Science, 52(4):621–639, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/187446. See addendum [684].

Cushing:1985:BRB

[427] James T. Cushing. Book review: Constructing Quarks: A Sociological History of Particle Physics by Andrew Pickering. Philosophy of Science, 52 (4):640–641, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187447.

Enfield:1985:BRB

[428] Patrick Enfield. Book review: Faraday to Einstein: Constructing Meaning in Scientific Theories by Nancy J. Nersessian. Philosophy of Science, 52 (4):641–642, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187448.

Healey:1985:BRB

[429] Richard Healey. Book review: Modern Logic and Quantum Mechanics by Rachel Wallace Garden. Philosophy of Science, 52(4):642–644, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187449.

Laudan:1985:BRB

[430] Rachel Laudan. Book review: The Dark Abyss of Time: The History of the Earth and the History of Nations from Hooke to Vico by Paolo Rossi; Lydia Cochrane. Philosophy of Science, 52(4):644–645, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187450.

Culver:1985:BRB

[431] David C. Culver. Book review: The Nature of Selection: Evolutionary Theory in Philosophical Focus by Elliott Sober. Philosophy of Science, 52 (4):645–646, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187451.

Leplin:1985:BRBb

[432] Jarrett Leplin. Book review: Is Science Progressive? by Ilkka Niiniluoto. Philosophy of Science, 52(4):646–648, December 1985. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187452.

Boler:1985:BRB

[433] John Boler. Book review: The Philosophy of Charles S. Peirce by Robert Almeder. Philosophy of Science, 52(4):648–649, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187453.

Anonymous:1985:BMd

[434] Anonymous. Back matter. Philosophy of Science, 52(4):651–660, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187454.

Anonymous:1985:VI

[435] Anonymous. Volume information. Philosophy of Science, 52(4):653–657, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187435.

Anonymous:1985:FMd

[436] Anonymous. Front matter. Philosophy of Science, 52(4):??, December 1985. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187436.

Mundy:1986:OAM

[437] Brent Mundy. Optical axiomatization of Minkowski space–time geometry. Philosophy of Science, 53(1):1–30, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187918.

Leplin:1986:MRS

[438] Jarrett Leplin. Methodological realism and scientific rationality. Philosophy of Science, 53(1):31–51, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187919.

Eells:1986:PCI

[439] Ellery Eells. Probabilistic causal interaction. Philosophy of Science, 53 (1):52–64, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187920.

Eagle:1986:BRB

[440] Morris N. Eagle. Book review: The Foundations of Psychoanalysis: A Philosophical Critique by Adolf Grünbaum. Philosophy of Science, 53(1): 65–88, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187921.

deMarchi:1986:MUP

[441] Neil de Marchi. Mill’s unrevised philosophy of economics: A comment on Hausman. Philosophy of Science, 53(1):89–100, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187922. See [112].

Zytkow:1986:WRD

[442] Jan M. Zytkow. What revisions does bootstrap testing need? Philosophy of Science, 53(1):101–109, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187923.

Gillies:1986:DPM

[443] Donald Gillies. In defense of the Popper-Miller argument. Philosophy of Science, 53(1):110–113, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187924. See response [556].

Nathan:1986:HSI

[444] Amos Nathan. How not to solve it. Philosophy of Science, 53(1):114– 119, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187925.

Moser:1986:PBR

[445] Paul K. Moser. Perception and belief: A regress problem. Philosophy of Science, 53(1):120–126, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187926.

Rosenberg:1986:WRP

[446] Alexander Rosenberg. What Rosenberg’s philosophy of economics is not. Philosophy of Science, 53(1):127–132, March 1986. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187927.

Forster:1986:CRB

[447] Malcolm R. Forster. Counterfactual reasoning in the Bell-epr paradox. Philosophy of Science, 53(1):133–144, March 1986. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187928.

Gunderson:1986:BRB

[448] Keith Gunderson. Book review: Matter and Consciousness: A Contemporary Introduction to the Philosophy of Mind by Paul M. Churchland. Philosophy of Science, 53(1):145–148, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187929.

Spielman:1986:BRB

[449] Stephen Spielman. Book review: Epistemology and Inference by Henry E. Kyburg. Philosophy of Science, 53(1):149–150, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187930.

Lambros:1986:BRB

[450] Charles H. Lambros. Book review: Quine and Analytic Philosophy by George D. Romanos. Philosophy of Science, 53(1):150–152, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187931.

Fuller:1986:BRB

[451] Steve Fuller. Book review: Understanding and Explanation: A Transcendental-Pragmatic Perspective by Karl-Otto Apel; Georgia Warnke. Philosophy of Science, 53(1):152–154, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http:/

/www.jstor.org/stable/187932.

Vallentyne:1986:BRB

[452] Peter Vallentyne. Book review: What Is a Law of Nature? by D. M. Armstrong. Philosophy of Science, 53(1):154–156, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187933.

Anonymous:1986:BMa

[453] Anonymous. Back matter. Philosophy of Science, 53(1):157–162, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187934.

Anonymous:1986:FMa

[454] Anonymous. Front matter. Philosophy of Science, 53(1):??, March 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187917.

Oddie:1986:PPP

[455] Graham Oddie. The poverty of the Popperian program for truthlikeness. Philosophy of Science, 53(2):163–178, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187690.

McCauley:1986:IRF

[456] Robert N. McCauley. Intertheoretic relations and the future of psychology. Philosophy of Science, 53(2):179–199, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187691.

Haimes:1986:HHM

[457] Y. Y. Haimes and A. Weiner. Hierarchical holographic modeling for conflict resolution. Philosophy of Science, 53(2):200–222, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187692.

Eells:1986:CCD

[458] Ellery Eells and Elliott Sober. Common causes and decision theory. Philosophy of Science, 53(2):223–245, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187693.

Pollock:1986:PP

[459] John L. Pollock. The paradox of the preface. Philosophy of Science, 53(2):246–258, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187694.

Derksen:1986:JPS

[460] A. A. Derksen. The justificational priority of science over the philosophy of science: Laudan’s science and hypothesis. Philosophy of Science, 53(2): 259–264, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187695.

Sperry:1986:MVM

[461] R. W. Sperry. Macro- versus micro-determinism. Philosophy of Science, 53(2):265–270, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187696.

Rapaport:1986:SET

[462] William J. Rapaport. Searle’s experiments with thought. Philosophy of Science, 53(2):271–279, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187697.

Redhead:1986:OEG

[463] M. L. G. Redhead and J. S. Steigerwald. Ontological economy and grand unified gauge theories. Philosophy of Science, 53(2):280–281, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187698.

Sorensen:1986:DIM

[464] Roy A. Sorensen. Did the intensity of my preferences double last night? Philosophy of Science, 53(2):282–285, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187699.

Weingard:1986:BRB

[465] Robert Weingard and Gerrit Smith. Book review: Foundations of Space– time Theories by Michael Friedman. Philosophy of Science, 53(2):286– 299, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187700.

Lavin:1986:BRB

[466] Michael Lavin. Book review: Hypothesis and Evidence in Psychoanalysis by Marshall Edelson. Philosophy of Science, 53(2):300–302, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187701.

Williams:1986:BRB

[467] Michael Williams. Book review: Science and Scepticism by John Watkins. Philosophy of Science, 53(2):302–305, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187702.

Bunge:1986:BRB

[468] Mario Bunge. Book review: Rational Thermodynamics by C. Truesdell. Philosophy of Science, 53(2):305–306, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187703.

Anonymous:1986:BR

[469] Anonymous. Books received. Philosophy of Science, 53(2):307–312, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187704.

Anonymous:1986:BMb

[470] Anonymous. Back matter. Philosophy of Science, 53(2):??, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187705.

Anonymous:1986:FMb

[471] Anonymous. Front matter. Philosophy of Science, 53(2):??, June 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187689.

Mulaik:1986:TSD

[472] Stanley A. Mulaik. Toward a synthesis of deterministic and probabilistic formulations of causal relations by the functional relation concept. Philosophy of Science, 53(3):313–332, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187672.

Robinson:1986:REQ

[473] Joseph D. Robinson. Reduction, explanation, and the quests of biological research. Philosophy of Science, 53(3):333–353, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187673.

Wachbroit:1986:PMO

[474] Robert Wachbroit. Progress: Metaphysical and otherwise. Philosophy of Science, 53(3):354–371, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187674.

Nier:1986:IHA

[475] Keith A. Nier. The importance of historical accuracy in philosophy of science: The case of Curd’s conception of Copernican rationality. Philosophy of Science, 53(3):372–394, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187675.

Ungar:1986:LTG

[476] Abraham Ungar. The Lorentz transformation group of the Special Theory of Relativity without Einstein’s isotropy convention. Philosophy of Science, 53(3):395–402, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187676.

Morrison:1986:QLI

[477] Margaret Morrison. Quantum logic and the invariance argument — a reply to Bell and Hallett. Philosophy of Science, 53(3):403–411, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187677. See [182].

Rosenberg:1986:FPP

[478] Alexander Rosenberg and Mary Williams. Fitness as primitive and propensity. Philosophy of Science, 53(3):412–418, September 1986. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187678.

Lugg:1986:ATM

[479] Andrew Lugg. An alternative to the traditional model? Laudan on disagreement and consensus in science. Philosophy of Science, 53(3):419–424, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187679.

Field:1986:BRB

[480] Hartry Field. Book review: Inquiry by Robert Stalnaker. Philosophy of Science, 53(3):425–448, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187680.

Suppes:1986:BRB

[481] Patrick Suppes. Book review: Science and Values: The Aims of Science and Their Role in Scientific Debate by Larry Laudan. Philosophy of Science, 53(3):449–451, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187681.

Zenzen:1986:BRB

[482] Michael J. Zenzen. Book review: Entropy in Relation to Incomplete Knowledge by K. G. Denbigh; J. S. Denbigh. Philosophy of Science, 53 (3):451–452, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187682.

Catton:1986:BRB

[483] Philip Catton. Book review: The Birth of History and Philosophy of Science: Kepler’s a Defence of Tycho against Ursus with Essays on Its Provenance and Significance by N. Jardine. Philosophy of Science, 53(3): 453–455, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187683.

Woodward:1986:BRB

[484] James Woodward. Book review: Data, Instruments, and Theory: A Dialectical Approach to Understanding Science by Robert John Ackermann. Philosophy of Science, 53(3):455–458, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187684.

MacKinnon:1986:BRB

[485] Edward MacKinnon. Book review: The Philosophy of Niels Bohr: The Framework of Complementarity by Henry J. Folse. Philosophy of Science, 53(3):458–459, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187685.

Binkley:1986:BRB

[486] Robert W. Binkley. Book review: Intending and Acting: Towards a Naturalized Action Theory by Myles Brand. Philosophy of Science, 53(3):

459–461, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187686.

Dupre:1986:BRB

[487] John Dupre. Book review: The Structure of Biological Science by Alexander Rosenberg. Philosophy of Science, 53(3):461–463, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187687.

Anonymous:1986:BMc

[488] Anonymous. Back matter. Philosophy of Science, 53(3):464–465, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187688.

Anonymous:1986:FMc

[489] Anonymous. Front matter. Philosophy of Science, 53(3):??, September 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187671.

Seidenfeld:1986:EU

[490] Teddy Seidenfeld. Entropy and uncertainty. Philosophy of Science, 53 (4):467–491, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187902.

Kincaid:1986:REI

[491] Harold Kincaid. Reduction, explanation, and individualism. Philosophy of Science, 53(4):492–513, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187903.

Tichy:1986:C

[492] Pavel Tichy. Constructions. Philosophy of Science, 53(4):514–534, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187904.

Clark:1986:PC

[493] Austen Clark. Psychofunctionalism and chauvinism. Philosophy of Science, 53(4):535–559, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187905.

Sterelny:1986:ID

[494] Kim Sterelny. The imagery debate. Philosophy of Science, 53(4):560–583, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187906.

Forge:1986:DAF

[495] John Forge. David Armstrong on functional laws. Philosophy of Science, 53(4):584–587, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187907.

Tait:1986:BRB

[496] W. W. Tait. Book review: Mathematics in Philosophy by Charles Parsons. Philosophy of Science, 53(4):588–606, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187908.

Rosenberg:1986:BRB

[497] Alexander Rosenberg. Book review: Vaulting Ambition: Sociobiology and the Quest for Human Nature by Philip Kitcher. Philosophy of Science, 53 (4):607–608, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187909.

Ruse:1986:BRB

[498] Michael Ruse. Book review: Evolution and Creation by Ernan McMullin. Philosophy of Science, 53(4):608–610, December 1986. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187910.

Watson:1986:BRB

[499] Richard A. Watson. Book review: The Great Devonian Controversy: The Shaping of Scientific Knowledge among Gentlemanly Specialists by Martin J. S. Rudwick. Philosophy of Science, 53(4):610–611, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187911.

Mayr:1986:BRB

[500] Ernst Mayr. Book review: Biophilosophical Implications of Inorganic and Organismic Evolution by Bernhard Rensch. Philosophy of Science, 53(4): 612, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187912.

Feverwerker:1986:BRB

[501] Elie Feverwerker. Book review: The Scientific Reinterpretation of Form by Norma E. Emerton. Philosophy of Science, 53(4):612–614, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187913.

Collier:1986:BRB

[502] John Collier. Book review: Evolution at a Crossroads: The New Biology and the New Philosophy of Science by David J. Depew; Bruce H. Weber. Philosophy of Science, 53(4):614–616, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187914.

VanGulick:1986:BRN

[503] Robert Van Gulick. Book review: The Nature of Psychological Explanation by Robert Cummins. Philosophy of Science, 53(4):616–618, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187915.

Anonymous:1986:BMd

[504] Anonymous. Back matter. Philosophy of Science, 53(4):619–629, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187916.

Anonymous:1986:VI

[505] Anonymous. Volume information. Philosophy of Science, 53(4):621–626, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187900.

Anonymous:1986:FMd

[506] Anonymous. Front matter. Philosophy of Science, 53(4):??, December 1986. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187901.

Skyrms:1987:DCP

[507] Brian Skyrms. Dynamic coherence and probability kinematics. Philosophy of Science, 54(1):1–20, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187470.

Woodward:1987:ITM

[508] James Woodward. On an information-theoretic model of explanation. Philosophy of Science, 54(1):21–44, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187471.

Otte:1987:ICC

[509] Richard Otte. Indeterminism, counterfactuals, and causation. Philosophy of Science, 54(1):45–62, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187472.

Smith:1987:IP

[510] Quentin Smith. Infinity and the past. Philosophy of Science, 54(1):63– 75, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187473.

Schoeman:1987:CIP

[511] Ferdinand Schoeman. Cohen on inductive probability and the law of evidence. Philosophy of Science, 54(1):76–91, March 1987. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187474.

Cohen:1987:ASF

[512] L. Jonathan Cohen. On analyzing the standards of forensic evidence: A reply to Schoeman. Philosophy of Science, 54(1):92–97, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187475.

Ruse:1987:SNP

[513] Michael Ruse. Is sociobiology a new paradigm? Philosophy of Science, 54(1):98–104, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187476.

Eells:1987:PCR

[514] Ellery Eells. Probabilistic causality: Reply to John Dupré. Philosophy of Science, 54(1):105–114, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187477. See [711].

Marras:1987:BRB

[515] Ausonio Marras. Book review: From Folk Psychology to Cognitive Science: The Case against Belief by Stephen Stich. Philosophy of Science, 54(1): 115–127, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187478.

Westfall:1987:BRB

[516] Richard S. Westfall. Book review: Leviathan and the Air-Pump: Hobbes, Boyle, and the Experimental Life by Steven Shapin; Simon Schaffer. Philosophy of Science, 54(1):128–130, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187479.

Wilson:1987:BRB

[517] Fred Wilson. Book review: The Concept of Physical Law by Norman Swartz. Philosophy of Science, 54(1):130–132, March 1987. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187480.

Healey:1987:BRB

[518] Richard Healey. Book review: Open Questions in Quantum Physics by Gino Tarozzi; Alwyn van der Merwe. Philosophy of Science, 54(1):132– 134, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187481.

Mundy:1987:BRB

[519] Brent Mundy. Book review: Physics as Natural Philosophy: Essays in Honor of Laszlo Tisza on His Seventy-Fifth Birthday by A. Shimony; H. Feshbach. Philosophy of Science, 54(1):134–135, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187482.

Kockelmans:1987:BRB

[520] Joseph J. Kockelmans. Book review: Die Entwicklung des Neuen Strukturalismus Seit 1973 by Wolfgang Stegmüller. Philosophy of Science, 54 (1):136–137, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187483.

Lugg:1987:BRB

[521] Andrew Lugg. Book review: The Limits of Scientific Reasoning by David Faust. Philosophy of Science, 54(1):137–138, March 1987. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187484.

Pitt:1987:BRB

[522] Joseph C. Pitt. Book review: Galileo and His Sources: The Heritage of the Collegio Romano in Galileo’s Science by William A. Wallace. Philosophy of Science, 54(1):138–140, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187485.

Morrison:1987:BRB

[523] Margaret Morrison. Book review: Descartes’ Philosophy of Science by Desmond Clarke. Philosophy of Science, 54(1):140–141, March 1987. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187486.

Arthur:1987:BRB

[524] Richard T. W. Arthur. Book review: Temporal Relations and Temporal Becoming: A Defense of a Russellian Theory of Time by L. Nathan Oaklander. Philosophy of Science, 54(1):142–143, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187487.

Anonymous:1987:BMa

[525] Anonymous. Back matter. Philosophy of Science, 54(1):144–145, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187488.

Anonymous:1987:FMa

[526] Anonymous. Front matter. Philosophy of Science, 54(1):??, March 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187469.

Recker:1987:CES

[527] Doren A. Recker. Causal efficacy: The structure of Darwin’s argument strategy in the origin of species. Philosophy of Science, 54(2):147–175, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187796.

Wicken:1987:EIS

[528] Jeffrey S. Wicken. Entropy and information: Suggestions for common language. Philosophy of Science, 54(2):176–193, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187797.

Laymon:1987:USD

[529] Ronald Laymon. Using Scott domains to explicate the notions of approximate and idealized data. Philosophy of Science, 54(2):194–221, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187798.

Rouse:1987:HPS

[530] Joseph Rouse. Husserlian phenomenology and scientific realism. Philosophy of Science, 54(2):222–232, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187799.

Eells:1987:COS

[531] Ellery Eells. Cartwright and Otte on Simpson’s Paradox. Philosophy of Science, 54(2):233–243, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187800.

Stone:1987:KLP

[532] Mark Stone. Kyburg, Levi, and Petersen. Philosophy of Science, 54(2): 244–255, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187801.

Misak:1987:PLA

[533] Cheryl Misak. Peirce, Levi, and the aims of inquiry. Philosophy of Science, 54(2):256–265, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187802.

Mayo:1987:HEC

[534] Deborah G. Mayo. How everyone can have a rare property: Response to Sober on frequency-dependent causation. Philosophy of Science, 54(2): 266–276, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187803. See

[366].

Hale:1987:SN

[535] Susan C. Hale and Michael D. Resnik. Science nominalized? Philosophy of Science, 54(2):277–280, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187804.

Horgan:1987:SNP

[536] Terence Horgan. Science nominalized properly. Philosophy of Science, 54(2):281–282, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187805.

Levin:1987:RDT

[537] Michael Levin. Rigid designators: Two applications. Philosophy of Science, 54(2):283–294, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187806.

VanBendegem:1987:ZPT

[538] Jean Paul Van Bendegem. Zeno’s paradoxes and the tile argument. Philosophy of Science, 54(2):295–302, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187807.

Bourgeois:1987:RFI

[539] Warren Bourgeois. On rejecting Foss’s image of van Fraassen. Philosophy of Science, 54(2):303–308, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187808.

Haugeland:1987:BRB

[540] John Haugeland. Book review: Computation and Cognition: Toward a Foundation for Cognitive Science by Zenon W. Pylyshyn. Philosophy of Science, 54(2):309–311, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187809.

Causey:1987:BRB

[541] Robert L. Causey. Book review: Explanation, Causation and Deduction by Fred Wilson. Philosophy of Science, 54(2):311–313, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187810.

Hooker:1987:BRBa

[542] C. A. Hooker. Book review: Popper and the Human Sciences by G. Currie; A. Musgrave. Philosophy of Science, 54(2):313–315, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187811.

Kitchener:1987:BRB

[543] Richard F. Kitchener. Book review: Psychogenèse et Histoire des Sciences by Jean Piaget; Rolando Garcia. Philosophy of Science, 54(2):315–317, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187812.

Zenzen:1987:BRB

[544] Michael Zenzen. Book review: Philosophy, Its History and Historiography by A. J. Holland. Philosophy of Science, 54(2):317–319, June 1987.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187813.

Anonymous:1987:BR

[545] Anonymous. Books received. Philosophy of Science, 54(2):320–324, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187814.

Anonymous:1987:BMb

[546] Anonymous. Back matter. Philosophy of Science, 54(2):325–326, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187815.

Anonymous:1987:FMb

[547] Anonymous. Front matter. Philosophy of Science, 54(2):??, June 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187795.

Norton:1987:LIO

[548] John Norton. The logical inconsistency of the old quantum theory of black body radiation. Philosophy of Science, 54(3):327–350, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187578.

Mitchell:1987:CUS

[549] Sandra D. Mitchell. Competing units of selection?: A case of symbiosis. Philosophy of Science, 54(3):351–367, September 1987. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187579. See correction [607].

Heelan:1987:HLP

[550] Patrick A. Heelan. Husserl’s later philosophy of natural science. Philosophy of Science, 54(3):368–390, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187580.

Nelson:1987:MMC

[551] R. J. Nelson. Machine models for cognitive science. Philosophy of Science, 54(3):391–408, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187581.

Achinstein:1987:SDM

[552] Peter Achinstein. Scientific discovery and Maxwell’s kinetic theory. Philosophy of Science, 54(3):409–434, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187582.

Kelly:1987:LD

[553] Kevin T. Kelly. The logic of discovery. Philosophy of Science, 54(3): 435–452, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187583.

Waters:1987:RLB

[554] C. Kenneth Waters. Relevance logic brings Hope to hypotheticodeductivism. Philosophy of Science, 54(3):453–464, September 1987. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187584.

Sober:1987:PLP

[555] Elliott Sober. Parsimony, likelihood, and the principle of the common cause. Philosophy of Science, 54(3):465–469, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187585.

Good:1987:RRG

[556] I. J. Good. A reinstatement, in response to Gillies, of Redhead’s argument in support of induction. Philosophy of Science, 54(3):470–472, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187586. See [443].

Abernethy:1987:GBC

[557] Brad Abernethy. Glymour on bootstrap confirmation of Ptolemaic theory. Philosophy of Science, 54(3):473–479, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187587.

Holsinger:1987:PSC

[558] Kent E. Holsinger. Pluralism and species concepts, or when must we agree with one another? Philosophy of Science, 54(3):480–485, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187588.

Michalos:1987:BRB

[559] Alex C. Michalos. Book review: The Politics and Rhetoric of Scientific Method by J. A. Schuster; R. R. Yeo. Philosophy of Science, 54(3):486, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187589.

Tweedale:1987:BRB

[560] Martin M. Tweedale. Book review: The Tradition of the Topics in the Middle Ages by Niels J. Green-Pedersen. Philosophy of Science, 54(3): 486–488, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187590.

McGrath:1987:BRB

[561] James H. McGrath. Book review: From Quarks to Quasars: Philosophical Problems of Modern Physics by Robert G. Colodny. Philosophy of Science, 54(3):488–489, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187591.

Nersessian:1987:BRB

[562] Nancy J. Nersessian. Book review: From Maxwell to Microphysics: Aspects of Electromagnetic Theory in the Last Quarter of the Nineteenth Century by Jed Z. Buchwald. Philosophy of Science, 54(3):489–490, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187592.

Anonymous:1987:BMc

[563] Anonymous. Back matter. Philosophy of Science, 54(3):??, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187593.

Anonymous:1987:FMc

[564] Anonymous. Front matter. Philosophy of Science, 54(3):??, September 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187577.

Irzik:1987:CMN

[565] Gurol Irzik and Eric Meyer. Causal modeling: New directions for statistical explanation. Philosophy of Science, 54(4):495–514, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187747.

Mayo:1987:MGS

[566] Deborah G. Mayo and Norman L. Gilinsky. Models of group selection. Philosophy of Science, 54(4):515–538, December 1987. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187748.

Davies:1987:HSM

[567] David Davies. Horwich on ‘semantic’ and ‘metaphysical’ realism. Philosophy of Science, 54(4):539–557, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187749.

Hellman:1987:EBC

[568] Geoffrey Hellman. EPR, Bell, and collapse: A route around “stochastic” hidden variables. Philosophy of Science, 54(4):558–576, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187750.

Albert:1987:QMA

[569] David Z. Albert. A quantum-mechanical automation. Philosophy of Science, 54(4):577–585, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187751.

Jeffrey:1987:IPJ

[570] Richard Jeffrey. Indefinite probability judgment: A reply to Levi. Philosophy of Science, 54(4):586–591, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187752.

Mayo:1987:CBR

[571] O. Mayo. Comments on “Randomization and the Design of Experiments” by P. Urbach. Philosophy of Science, 54(4):592–596, December 1987. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187753. See [384].

Fetzer:1987:BRB

[572] James H. Fetzer. Book review: Scientific Explanation and the Causal Structure of the World by Wesley Salmon. Philosophy of Science, 54(4): 597–610, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187754.

Hooker:1987:BRBb

[573] Cliff Hooker. Book review: The Quantum and beyond by W. M. Honig. Philosophy of Science, 54(4):611–612, December 1987. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187755.

Giaquinto:1987:BRB

[574] M. Giaquinto. Book review: The Rationality of Induction by D. C. Stove. Philosophy of Science, 54(4):612–615, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187756.

Shea:1987:BRB

[575] William R. Shea. Book review: Il Neopositivismo Logico by Francesco Barone. Philosophy of Science, 54(4):615–616, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187757.

Anonymous:1987:BMd

[576] Anonymous. Back matter. Philosophy of Science, 54(4):617–631, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187758.

Anonymous:1987:VI

[577] Anonymous. Volume information. Philosophy of Science, 54(4):619–624, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187745.

Anonymous:1987:FMd

[578] Anonymous. Front matter. Philosophy of Science, 54(4):??, December 1987. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187746.

Kitcher:1988:MCT

[579] Patricia Kitcher. Marr’s computational theory of vision. Philosophy of Science, 55(1):1–24, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187817.

Hintikka:1988:IT

[580] Jaakko Hintikka. On the incommensurability of theories. Philosophy of Science, 55(1):25–38, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187818.

Smith:1988:UBU

[581] Quentin Smith. The uncaused beginning of the universe. Philosophy of Science, 55(1):39–57, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187819.

Sherry:1988:ZMP

[582] David M. Sherry. Zeno’s metrical paradox revisited. Philosophy of Science, 55(1):58–73, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187820.

Thompson:1988:SPW

[583] Paul Thompson. Some punctuationists are wrong about the modern synthesis. Philosophy of Science, 55(1):74–86, March 1988. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187821.

Pitt:1988:GRE

[584] Joseph C. Pitt. Galileo, rationality and explanation. Philosophy of Science, 55(1):87–103, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187822.

Doppelt:1988:PRA

[585] Gerald Doppelt. The philosophical requirements for an adequate conception of scientific rationality. Philosophy of Science, 55(1):104–133, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187823.

Shapere:1988:DC

[586] Dudley Shapere. Doppelt crossed. Philosophy of Science, 55(1):134–140, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187824.

Temple:1988:CTW

[587] Dennis Temple. The contrast theory of why-questions. Philosophy of Science, 55(1):141–151, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187825.

Clarke:1988:BRB

[588] Murray Clarke. Book review: Naturalizing Epistemology by Hilary Kornblith. Philosophy of Science, 55(1):152–153, March 1988. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187826.

Jones:1988:BRB

[589] Roger Jones. Book review: Light in Einstein’s Universe: The Role of Energy in Cosmology and Relativity by S. J. Prokhovnik. Philosophy of Science, 55(1):153–155, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187827.

Teller:1988:BRB

[590] Paul Teller. Book review: The Shaky Game: Einstein, Realism, and the Quantum Theory by Arthur Fine. Philosophy of Science, 55(1):155– 156, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187828.

Anonymous:1988:BMa

[591] Anonymous. Back matter. Philosophy of Science, 55(1):157–166, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187829.

Anonymous:1988:FMa

[592] Anonymous. Front matter. Philosophy of Science, 55(1):??, March 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187816.

Churchland:1988:PPT

[593] Paul M. Churchland. Perceptual plasticity and theoretical neutrality: A reply to Jerry Fodor. Philosophy of Science, 55(2):167–187, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187956.

Fodor:1988:RCP

[594] Jerry A. Fodor. A reply to Churchland’s “perceptual plasticity and theoretical neutrality”. Philosophy of Science, 55(2):188–198, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187957.

Levin:1988:MRR

[595] Janet Levin. Must reasons be rational? Philosophy of Science, 55(2):199– 217, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187958.

Feldman:1988:RRN

[596] Richard Feldman. Rationality, reliability, and natural selection. Philosophy of Science, 55(2):218–227, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187959.

Sober:1988:LC

[597] Elliott Sober. Likelihood and convergence. Philosophy of Science, 55(2): 228–237, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187960.

deQueiroz:1988:SDR

[598] Kevin de Queiroz. Systematics and the Darwinian revolution. Philosophy of Science, 55(2):238–259, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187961.

Earman:1988:WRD

[599] John Earman and Clark Glymour. What revisions does bootstrap testing need? A reply. Philosophy of Science, 55(2):260–264, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187962.

Edidin:1988:RCR

[600] Aron Edidin. From relative confirmation to real confirmation. Philosophy of Science, 55(2):265–271, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187963.

Duran:1988:CRE

[601] Jane Duran. Causal reference and epistemic justification. Philosophy of Science, 55(2):272–279, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187964.

Catton:1988:UES

[602] Philip Catton and Graham Solomon. Uniqueness of embeddings and space–time relationalism. Philosophy of Science, 55(2):280–291, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187965.

Wylie:1988:CNB

[603] Alison Wylie. Critical notice: Testing Scientific Theories, John Earman (Ed.): Explaining Confirmation Practice. Philosophy of Science, 55(2): 292–303, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187966.

Bunn:1988:BRBa

[604] Robert Bunn. Book review: Time, Creation, and the Continuum: Theories in Antiquity and the Early Middle Ages by Richard Sorabji. Philosophy of Science, 55(2):304–306, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187967.

Hacking:1988:BRB

[605] Ian Hacking. Book review: The Neglect of Experiment by Allan Franklin. Philosophy of Science, 55(2):306–308, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187968.

Anonymous:1988:BR

[606] Anonymous. Books received. Philosophy of Science, 55(2):309–313, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187969.

Anonymous:1988:CCU

[607] Anonymous. Correction: Competing units of selection?: A case of symbiosis. Philosophy of Science, 55(2):314, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187970. See [549].

Anonymous:1988:BMb

[608] Anonymous. Back matter. Philosophy of Science, 55(2):315–321, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187971.

Anonymous:1988:FMb

[609] Anonymous. Front matter. Philosophy of Science, 55(2):??, June 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187955.

Splitter:1988:SI

[610] Laurance J. Splitter. Species and identity. Philosophy of Science, 55(3): 323–348, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187652.

Bitbol:1988:CMT

[611] M. Bitbol. The concept of measurement and time symmetry in quantum mechanics. Philosophy of Science, 55(3):349–375, September 1988. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187653.

Rota:1988:SSP

[612] Gian-Carlo Rota, David H. Sharp, and Robert Sokolowski. Syntax, semantics, and the problem of the identity of mathematical objects. Philosophy of Science, 55(3):376–386, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187654.

Svetlichny:1988:DBI

[613] George Svetlichny, Michael Redhead, Harvey Brown, and Jeremy Butterfield. Do the Bell inequalities require the existence of joint probability distributions? Philosophy of Science, 55(3):387–401, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187655.

Landsberg:1988:CPS

[614] P. T. Landsberg and J. Wise. Components of probabilistic support: The two-proposition case. Philosophy of Science, 55(3):402–414, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187656.

Couvalis:1988:FLB

[615] Spyridon George Couvalis. Feyerabend and Laymon on Brownian motion. Philosophy of Science, 55(3):415–421, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187657.

Neander:1988:WDN

[616] Karen Neander. What does natural selection explain? Correction to Sober. Philosophy of Science, 55(3):422–426, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187658.

Ereshefsky:1988:AIR

[617] Marc Ereshefsky. Axiomatics and individuality: A reply to Williams’ “Species Are Individuals”. Philosophy of Science, 55(3):427–434, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187659. See [421].

Baigrie:1988:SRS

[618] Brian S. Baigrie. Siegel on the rationality of science. Philosophy of Science, 55(3):435–441, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 187660. See response [619].

Siegel:1988:RAR

[619] Harvey Siegel. Rationality and anemia (response to Baigrie). Philosophy of Science, 55(3):442–447, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187661. See [618].

Noorbaloochi:1988:HJA

[620] S. Noorbaloochi. On Harold Jeffreys’ axioms. Philosophy of Science, 55 (3):448–452, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187662.

Eells:1988:QSI

[621] Ellery Eells. Quentin Smith on infinity and the past. Philosophy of Science, 55(3):453–455, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187663.

Dieks:1988:SRF

[622] D. Dieks. Special Relativity and the flow of time. Philosophy of Science, 55 (3):456–460, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187664.

Bunn:1988:BRBb

[623] Robert Bunn. Book review: Cantorian Set Theory and Limitation of Size by Michael Hallett. Philosophy of Science, 55(3):461–478, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187665.

Freed:1988:BRB

[624] Bruce Freed. Book review: Epistemology and Cognition by Alvin I. Goldman. Philosophy of Science, 55(3):479–480, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187666.

Garrett:1988:BRB

[625] Don Garrett. Book review: Spinoza and the Sciences by Marjorie Grene; Debra Nails. Philosophy of Science, 55(3):480–482, September 1988. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187667.

Hon:1988:BRB

[626] Giora Hon. Book review: Observation, Experiment, and Hypothesis in Modern Physical Science by Peter Achinstein; Owen Hannaway. Philosophy of Science, 55(3):482–486, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187668.

Lugg:1988:BRB

[627] Andrew Lugg. Book review: Wittgenstein: Rules, Grammar and Necessity by G. P. Baker; P. M. S. Hacker. Philosophy of Science, 55(3):486–487, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187669.

Anonymous:1988:BMc

[628] Anonymous. Back matter. Philosophy of Science, 55(3):488–492, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187670.

Anonymous:1988:FMc

[629] Anonymous. Front matter. Philosophy of Science, 55(3):??, September 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187651.

Leplin:1988:EU

[630] Jarrett Leplin. Is essentialism unscientific? Philosophy of Science, 55 (4):493–510, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187525.

Foss:1988:PVF

[631] Jeff Foss. The percept and vector function theories of the brain. Philosophy of Science, 55(4):511–537, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187526.

Forster:1988:SPC

[632] Malcolm R. Forster. Sober’s Principle of Common Cause and the problem of comparing incomplete hypotheses. Philosophy of Science, 55(4):538– 559, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187527.

Weirich:1988:HMT

[633] Paul Weirich. Hierarchical maximization of two kinds of expected utility. Philosophy of Science, 55(4):560–582, December 1988. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187528.

Goldstick:1988:T

[634] D. Goldstick and B. O’Neill. “truer”. Philosophy of Science, 55(4):583– 597, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187529.

Vallentyne:1988:EL

[635] Peter Vallentyne. Explicating lawhood. Philosophy of Science, 55(4): 598–613, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187530.

Bigelow:1988:F

[636] John Bigelow, Brian Ellis, and Robert Pargetter. Forces. Philosophy of Science, 55(4):614–630, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187531.

Kronz:1988:ECS

[637] Frederick M. Kronz. EPR: The correlations are still a mystery. Philosophy of Science, 55(4):631–639, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187532.

Maxwell:1988:PSR

[638] Nicholas Maxwell. Are probabilism and Special Relativity compatible? Philosophy of Science, 55(4):640–645, December 1988. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187533.

Levi:1988:BRB

[639] Isaac Levi. Book review: Probabilistic Metaphysics by Patrick Suppes. Philosophy of Science, 55(4):646–652, December 1988. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187534.

Dennett:1988:BRB

[640] Daniel C. Dennett. Book review: The Disappearance of Introspection by William Lyons. Philosophy of Science, 55(4):653–654, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187535.

Sterelny:1988:BRB

[641] Kim Sterelny. Book review: Knowing Who by Stephen E. Boër; William G. Lycan. Philosophy of Science, 55(4):654–656, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187536.

Anonymous:1988:BMd

[642] Anonymous. Back matter. Philosophy of Science, 55(4):657–669, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187537.

Anonymous:1988:VI

[643] Anonymous. Volume information. Philosophy of Science, 55(4):659–664, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187523.

Anonymous:1988:FMd

[644] Anonymous. Front matter. Philosophy of Science, 55(4):??, December 1988. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187524.

Martin:1989:ICP

[645] Jane Roland Martin. Ideological critiques and the philosophy of science. Philosophy of Science, 56(1):1–22, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187539.

Tuomela:1989:MSE

[646] Raimo Tuomela. Methodological solipsism and explanation in psychology. Philosophy of Science, 56(1):23–47, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187540.

Kaplan:1989:BBB

[647] Mark Kaplan. Bayesianism without the black box. Philosophy of Science, 56(1):48–69, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187541.

Sobel:1989:PTC

[648] Jordan Howard Sobel. Partition-theorems for causal decision theories. Philosophy of Science, 56(1):70–93, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187542.

Osherson:1989:ICC

[649] Daniel N. Osherson and Scott Weinstein. Identifiable collections of countable structures. Philosophy of Science, 56(1):94–105, March 1989. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187543.

Darden:1989:STT

[650] Lindley Darden and Joseph A. Cain. Selection type theories. Philosophy of Science, 56(1):106–129, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187544.

Schroeder-Heister:1989:RRC

[651] Peter Schroeder-Heister and Frank Schaefer. Reduction, representation and commensurability of theories. Philosophy of Science, 56(1):130–157, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187545. See note [652].

Pearce:1989:TRC

[652] David Pearce. Translation, reduction and commensurability: A note on Schroeder-Heister and Schaefer. Philosophy of Science, 56(1):158–164, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/187546. See [651].

Postow:1989:FMP

[653] B. C. Postow. The FPL model and practical inference. Philosophy of Science, 56(1):165–170, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187547.

Stich:1989:BRB

[654] Stephen P. Stich. Book review: Minimal Rationality by Christopher Cherniak. Philosophy of Science, 56(1):171–173, March 1989. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187548.

Sayre-McCord:1989:BRB

[655] Geoffrey Sayre-McCord. Book review: The Natural Philosophy of Leibniz by K. Okruhlik; J. Brown. Philosophy of Science, 56(1):173–174, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187549.

Beatty:1989:BRB

[656] John Beatty. Book review: A History of Embryology by T. J. Horder; J. A. Witkowski; C. C. Wylie. Philosophy of Science, 56(1):174–177, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187550.

Hodes:1989:BRB

[657] Harold T. Hodes. Book review: Intensional Mathematics by Stewart Shapiro. Philosophy of Science, 56(1):177–178, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187551.

Anonymous:1989:BMa

[658] Anonymous. Back matter. Philosophy of Science, 56(1):179–184, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187552.

Anonymous:1989:FMa

[659] Anonymous. Front matter. Philosophy of Science, 56(1):??, March 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187538.

Kelly:1989:CTN

[660] Kevin T. Kelly and Clark Glymour. Convergence to the truth and nothing but the truth. Philosophy of Science, 56(2):185–220, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187870.

Laudan:1989:DDM

[661] Rachel Laudan and Larry Laudan. Dominance and the disunity of method: Solving the problems of innovation and consensus. Philosophy of Science, 56(2):221–237, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187871.

Teller:1989:IR

[662] Paul Teller. Infinite renormalization. Philosophy of Science, 56(2):238– 257, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187872.

McClamrock:1989:HTL

[663] Ron McClamrock. Holism without tears: Local and global effects in cognitive processes. Philosophy of Science, 56(2):258–274, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187873.

Sober:1989:IEA

[664] Elliott Sober. Independent evidence about a common cause. Philosophy of Science, 56(2):275–287, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187874.

Millikan:1989:DPF

[665] Ruth Garrett Millikan. In defense of proper functions. Philosophy of Science, 56(2):288–302, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187875.

Baker:1989:II

[666] Lynne Rudder Baker. Instrumental intentionality. Philosophy of Science, 56(2):303–316, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187876.

Graves:1989:TET

[667] Paul R. Graves. The total evidence theorem for probability kinematics. Philosophy of Science, 56(2):317–324, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187877.

Leeds:1989:MZG

[668] Stephen Leeds. Malament and Zabell on Gibbs phase averaging. Philosophy of Science, 56(2):325–340, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187878. See [39].

Rosenberg:1989:RVS

[669] Alexander Rosenberg. Russell versus Steiner on physics and causality. Philosophy of Science, 56(2):341–347, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187879.

Bar-Hillel:1989:HSP

[670] Maya Bar-Hillel. How to solve probability teasers. Philosophy of Science, 56(2):348–358, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187880.

Egan:1989:BRB

[671] M. Frances Egan. Book review: Thoughts: An Essay on Content by Christopher Peacocke. Philosophy of Science, 56(2):359–360, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187881.

Mundy:1989:BRB

[672] Brent Mundy. Book review: Methodology, Epistemology, and Philosophy: Essays in Honour of Wolfgang Stegmüller on the Occasion of His 60th Birthday by Carl G. Hempel; Hilary Putnam; Wilhelm K. Essler. Philosophy of Science, 56(2):361–362, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187882.

Anonymous:1989:BR

[673] Anonymous. Books received. Philosophy of Science, 56(2):363–366, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187883.

Anonymous:1989:BMb

[674] Anonymous. Back matter. Philosophy of Science, 56(2):367–372, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187884.

Anonymous:1989:FMb

[675] Anonymous. Front matter. Philosophy of Science, 56(2):??, June 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187869.

Grunbaum:1989:PPC

[676] Adolf Grünbaum. The pseudo-problem of Creation in physical cosmology. Philosophy of Science, 56(3):373–394, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187991.

Lloyd:1989:SAD

[677] Elisabeth A. Lloyd. A structural approach to defining units of selection. Philosophy of Science, 56(3):395–418, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187992.

Shapere:1989:ECS

[678] Dudley Shapere. Evolution and continuity in scientific change. Philosophy of Science, 56(3):419–437, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187993.

Blachowicz:1989:DAI

[679] James Blachowicz. Discovery and ampliative inference. Philosophy of Science, 56(3):438–462, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187994.

Silvers:1989:RCI

[680] Stuart Silvers. Representational capacity, intentional ascription, and the slippery slope. Philosophy of Science, 56(3):463–473, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187995.

Sapire:1989:DRP

[681] David Sapire. Determinism, the remote past, and the causal or determinational structure of the universe. Philosophy of Science, 56(3):474–483, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187996.

Shanahan:1989:BCN

[682] Timothy Shanahan. Beatty on chance and natural selection. Philosophy of Science, 56(3):484–489, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187997.

Townsend:1989:PDS

[683] Burke Townsend. Partly deductive support in the Popper-Miller argument. Philosophy of Science, 56(3):490–496, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187998.

Solomon:1989:ADF

[684] Graham Solomon. An addendum to Demopoulos and Friedman (1985). Philosophy of Science, 56(3):497–501, September 1989. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187999. See [426].

Wilson:1989:CNJ

[685] Mark Wilson. Critical notice: John Earman’s A Primer on Determinism. Philosophy of Science, 56(3):502–532, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188000.

Stabler:1989:BRB

[686] Edward P. Stabler, Jr. Book reviews: Knowledge of Language: Its Nature, Origin, and Use by Noam Chomsky; Language and Problems of Knowledge: The Managua Lectures by Noam Chomsky. Philosophy of Science, 56 (3):533–536, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/188001.

Gale:1989:BRB

[687] George Gale. Book review: The Advancement of Science, and Its Burdens by Gerald Holton. Philosophy of Science, 56(3):536–537, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/188002.

Brittan:1989:BRB

[688] Gordon G. Brittan, Jr. Book review: Foundations of Objective Knowledge: The Relations of Popper’s Theory of Knowledge to That of Kant by Sergio L. de C. Fernandes. Philosophy of Science, 56(3):537–539, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/188003.

Churchland:1989:BRB

[689] Patricia Smith Churchland. Book review: Memory and Brain by Larry R. Squire. Philosophy of Science, 56(3):539–540, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188004.

Dennett:1989:BRB

[690] Daniel C. Dennett. Book review: Darwin and the Emergence of Evolutionary Theories of Mind and Behavior by Robert J. Richards. Philosophy of Science, 56(3):540–543, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/188005.

Thornton:1989:BRB

[691] Mark Thornton. Book review: Elbow Room: The Varieties of Free Will worth Wanting by Daniel C. Dennett. Philosophy of Science, 56(3):543– 544, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/188006.

Dretske:1989:BRB

[692] Fred Dretske. Book review: Observation and Objectivity by Harold I. Brown. Philosophy of Science, 56(3):544–547, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/188007.

Anonymous:1989:BMc

[693] Anonymous. Back matter. Philosophy of Science, 56(3):548–553, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/188008.

Anonymous:1989:FMc

[694] Anonymous. Front matter. Philosophy of Science, 56(3):??, September 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187990.

Hacking:1989:ERC

[695] Ian Hacking. Extragalactic reality: The case of gravitational lensing. Philosophy of Science, 56(4):555–581, December 1989. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187781.

Mundy:1989:QRT

[696] Brent Mundy. On quantitative relationist theories. Philosophy of Science, 56(4):582–600, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187782.

Scharf:1989:QMP

[697] David C. Scharf. Quantum measurement and the program for the unity of science. Philosophy of Science, 56(4):601–623, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187783.

Rabinowicz:1989:SRO

[698] Wlodzimierz Rabinowicz. Stable and retrievable options. Philosophy of Science, 56(4):624–641, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187784.

Menzies:1989:PCC

[699] Peter Menzies. Probabilistic causation and causal processes: A critique of Lewis. Philosophy of Science, 56(4):642–663, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/187785.

Egan:1989:WWS

[700] M. Frances Egan. What’s wrong with the syntactic theory of mind. Philosophy of Science, 56(4):664–674, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187786.

Howson:1989:ROP

[701] Colin Howson. On a recent objection to Popper and Miller’s “disproof” of probabilistic induction. Philosophy of Science, 56(4):675–680, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187787.

Eriksen:1989:CPL

[702] Leif Eriksen. Confirmation, paradox, and logic. Philosophy of Science, 56 (4):681–687, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/187788.

Clifton:1989:SRC

[703] Robert K. Clifton. Some recent controversy over the possibility of experimentally determining isotropy in the speed of light. Philosophy of Science, 56(4):688–696, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

187789.

Gupta:1989:BRB

[704] Anil Gupta. Book review: The Liar: An Essay on Truth and Circularity by Jon Barwise; John Etchemendy. Philosophy of Science, 56(4):697–709, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187790.

Batterman:1989:BRB

[705] Robert Batterman. Book review: Time, The Physical Magnitude by Olivier Costa de Beauregard. Philosophy of Science, 56(4):710–712, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187791.

Stairs:1989:BRB

[706] Allen Stairs. Book review: Particles and Paradoxes: The Limits of Quantum Logic by Peter Gibbins. Philosophy of Science, 56(4):712–714, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187792.

DiSalle:1989:BRB

[707] Robert DiSalle. Book review: Philosophy and Spacetime Physics by Lawrence Sklar. Philosophy of Science, 56(4):714–717, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/187793.

Anonymous:1989:BMd

[708] Anonymous. Back matter. Philosophy of Science, 56(4):719–732, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187794.

Anonymous:1989:VI

[709] Anonymous. Volume information. Philosophy of Science, 56(4):721–726, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/187779.

Anonymous:1989:FMd

[710] Anonymous. Front matter. Philosophy of Science, 56(4):??, December 1989. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/187780.

Dupre:1990:PCR

[711] John Dupré. Probabilistic causality: A rejoinder to Ellery Eells. Philosophy of Science, 57(4):690–698, December 1990. CODEN PHSCA6. ISSN REFERENCES 113

0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/187770. See [514].